### v2

# Exergetic Port-Hamiltonian Systems Modeling Language

Abstract
Mathematical modeling of real-world physical systems requires the consistent combination of a multitude of physical laws and phenomenological models. This challenging task can be greatly simplified by hierarchically decomposing systems into ultimately simple components. Moreover, the use of diagrams for expressing the decomposition helps make the process more intuitive and facilitates communication, even with non-experts. As an important requirement, models have to respect fundamental physical laws such as the first and the second law of thermodynamics. While some existing modeling frameworks make such guarantees based on structural properties of their models, they lack a formal graphical syntax. We present a compositional and thermodynamically consistent modeling language with a graphical syntax. In terms of its semantics, we essentially endow port-Hamiltonian systems with additional structural properties and a fixed physical interpretation, ensuring thermodynamic consistency in a manner closely related to the metriplectic or GENERIC formalism. While port-Hamiltonian systems are inspired by graphical modeling with bond graphs, neither the link between the two, nor bond graphs themselves, can be easily formalized. In contrast, our syntax is based on a refinement of the well-studied operad of undirected wiring diagrams. By combining a compositional, graphical syntax with an energy-based, thermodynamic approach, the presented modeling language simplifies the understanding, reuse, and modification of complex physical models.
keywords:
bond graphs , compositionality , GENERIC , metriplectic structure , nonequilibrium thermodynamics , undirected wiring diagrams
journal: arXiv
1 Introduction
In this introduction we mainly want to answer the following questions:
  1. 1.
    What are Exergetic Port-Hamiltonian Systems (EPHS)?
  2. 2.
    How do they enable compositional modeling?
  3. 3.
    Why is this practically relevant?
In addition, we briefly summarize key ideas of nonequilibrium thermodynamics, and we discuss related work that inspired our developments. The introduction finishes with an outline of the main part of the paper.
1.1 What?
Exergetic Port-Hamiltonian Systems (EPHS) provide a compositional and thermodynamically consistent language for expressing mathematical models of multiphysical systems at macroscopic scales. Such systems may combine
  1. 1.
    classical mechanics
    (e.g. spring-mass systems, multibody dynamics, fluid mechanics),
  2. 2.
    classical electromagnetism
    (e.g. LC circuits, electromagnetic wave propagation)
  3. 3.
    and irreversible processes with local thermodynamic equilibrium
    (e.g. friction/viscosity, electrical resistance/conduction, thermal conduction).
The EPHS language is termed compositional because it enables a fully modular and hierarchical approach to model specification through its simple, graphical syntax, as elaborated in the next subsection.
Flattening any hierarchical nesting reveals that every EPHS model is fundamentally defined by a power-preserving interconnection of primitive subsystems. There are three main classes of primitive EPHS:
  1. 1.
    Systems that represent energy storage are characterized by a function which yields the stored energy for each state of the primitive system.
  2. 2.
    Systems that represent reversible energy exchange are characterized by a so-called Dirac structure [1], as known e.g. from (constrained) mechanical systems.
  3. 3.
    Systems that represent irreversible energy exchange are characterized by another type of structure that we call Onsager structure.
Since both classical mechanics and electromagnetism can be expressed using Dirac structures, item 2. corresponds to (a) and (b) above, while item 3. corresponds to (c). Structural properties of the primitive systems and their interconnection guarantee that all models respect the first and the second law of thermodynamics, as well as Onsager’s reciprocal relations.
Inspired by the contribution in [2], previous work [3] has explored EPHS as a systematic approach to thermodynamic modeling based on port-Hamiltonian systems (see, e.g., [4]). By interpreting the ‘Hamiltonian’ as an exergy storage function, a stricter version of port-Hamiltonian structure can be formulated, ensuring a thermodynamically consistent reversible-irreversible splitting of the dynamics, as known from the metriplectic or GENERIC formalism (see, e.g., [5, 6]). The exergy storage function represents the theoretically ‘useful’ amount of energy, i.e., the maximum amount of work that can be extracted from a thermodynamic system using an ideal device, such as a Carnot engine. This ideal device has access not only to all forms of energy stored in the system but also to a (reference) environment, primarily characterized by a heat bath with constant temperature.
The present work formalizes EPHS as a compositional modeling language with a simple, graphical syntax inspired by bond graphs (see, e.g., [7, 8]). This approach enables the construction of complex models by composing them from simpler subsystems, which can themselves be constructed from more basic subsystems, and so on. We draw on and extend ideas form (applied) category theory (see, e.g., [9, 10]) to propose a simple mathematical framework suitable for computer implementation. In principle, the graphical syntax and the associated compositional framework could also be adopted without the parts that ensure thermodynamic consistency. While this would relax certain conditions imposed on the primitive systems, it would also weaken the correspondence between the remaining structure and its physical interpretation.
1.2 How?
Figure 1: Graphical display of an interconnection pattern with different syntactic elements shown in different colors. The outer box shown in purple represents the interface of the resulting composite system. The three protruding lines represent its ports called $\mathtt{f.p}$, $\mathtt{f.m}$ and $\mathtt{b_{k}}$. The two inner boxes shown in red represent (the interface of) the subsystems called $\mathtt{ke}$ (storage of kinetic energy) and $\mathtt{sa}$ (self-advection of kinetic energy). Both have ports called $\mathtt{p}$ (momentum) and $\mathtt{m}$ (mass). Box $\mathtt{sa}$ has another port called $\mathtt{b_{k}}$ (advection of kinetic energy across the boundary of the spatial domain). The common prefix $\mathtt{f}$ (fluid) of the outer ports $\mathtt{f.p}$ (fluid momentum) and $\mathtt{f.m}$ (fluid mass) combines them into a multiport. The black dots represent junctions where power is exchanged among the connected ports. The dashed line indicates that only information about the state (fluid mass) is exchanged via the outer port $\mathtt{f.m}$, with no energy transfer occurring.
Expressions in the graphical syntax are called interconnection patterns, since such a pattern specifies how a composite system is formed through a power-preserving interconnection of subsystems. Figure 1 shows an example from fluid dynamics, see [11].
The syntax deals with systems only in terms of their interfaces, i.e. their exposed ports. Each port has two attributes: The first attribute gives the physical quantity that can be exchanged via the port together with the corresponding state space. To rule out ill-defined connections, only ports with the same quantity can be connected. The second attribute is of Boolean type and indicates whether power can be exchanged over the port or merely information about system state, as indicated graphically by a solid or a dashed line, respectively.
We think of an interconnection pattern as a function. For each inner box, it takes as input a system with matching interface and it outputs a system whose interface matches the outer box. For instance, the pattern shown in Figure 1 represents a function that takes two subsystems, whose interfaces are determined by the ports of the boxes $\mathtt{ke}$ and $\mathtt{sa}$, and it combines them into a composite system, whose interface is given by the ports of the outer box. We draw inner boxes filled with a color to indicate the nature of the subsystems that we consider as inputs. In Figure 1, box $\mathtt{ke}$ has a blue filling, as it represents energy storage (the fluid’s kinetic energy). Box $\mathtt{sa}$ has a green filling, as it represents reversible energy exchange (advection of momentum and mass). Written on top of the outer box, kin is an identifier akin to a function name and the text in parentheses is just a short description. The multiport $\mathtt{f}$ facilitates integrating the system into a more complex model without drawing parallel connections.
(a) Kinetic energy system
(b) Fluid system
(c) EMHD system
Figure 2: Hierarchical specification of an electro-magneto hydrodynamics (EMHD) model [11]: The fluid model f shown in (b) includes the kinetic energy system kin depicted in (a), as well as four additional models not detailed here: an internal energy system int, a thermal conduction model th, and volume and shear viscosity models vol and shr. The electro-magneto hydrodynamics model emhd shown in (c) includes the fluid model f depicted in (b), as well as three additional models not detailed here: an electromagnetic system em, an electric conduction model el, and a model for the electro-kinetic coupling ekc.
Figure 2 shows the hierarchical specification of a plasma model. For simplicity, a spatial domain with isolated boundary is assumed such that only the in-domain interconnection of the different models needs to be dealt with. A red filling indicates that the respective input is a primitive system modeling an irreversible process. A yellow filling signifies that the input is a nested composite system, specified in terms of its own pattern. Recursively substituting into all yellow boxes the respective patterns would yield an equivalent model specification without the hierarchical decomposition. The composition of patterns through substitution is uniquely defined whenever interfaces match (possibly up to a given renaming of the ports). No additional data or handling of ‘edge cases’ is required. One can hence easily add levels of abstraction by refactoring subsystems into reusable parts. In the fluid model, one could for instance factor out the kinetic energy system filling box $\mathtt{kin}$ and the internal energy system filling box $\mathtt{int}$ into a reusable model of an ideal compressible fluid, see [11].
1.3 Why?
EPHS provide a formal and intuitive modeling language suited for humans and software tools alike. As shown above, increasingly complex models are assembled from simpler and ultimately primitive parts. This fosters the reusability of models and encourages the conceptual separation of different levels of detail. Beyond the modeling process, this can be beneficial when the language is used in interdisciplinary communication and education. Some modeling errors can be completely avoided due to the built-in thermodynamic consistency and the typed interfaces (given by ports with fixed physical quantities). Since interconnection patterns are essentially energy/exergy flow diagrams, EPHS models are particularly useful for thermodynamic analysis and optimization. The compositional nature of EPHS could enable a scalable approach to model transformations, such as discretizations. It could also enable a practical integration of first-principles-based thermodynamic modeling and scientific machine learning.
1.4 Nonequilibrium thermodynamics in a nutshell
Most physical systems are characterized by a huge number of degrees of freedom but keeping track of them is usually neither possible nor useful. Nonequilibrium thermodynamics aims to provide a framework for coarse-grained modeling, meaning that fast dynamics at small scales are not resolved, and a quantity called entropy is introduced to account for the missing information, see, e.g., [6]. Coarse-graining is possible because, despite fluctuations at the microscopic scale, one can eventually observe non-fluctuating properties at a more macroscopic scale; at least this is the case for isolated systems. For instance, temperature is a property that only appears at more macroscopic scales. Thermodynamic properties are classified as extensive when they are proportional to the size of the system (e.g. entropy, volume, mass) or as intensive when they do not depend on the size of the system (e.g. temperature, pressure, chemical potential), see, e.g., [12]
A system is said to be in local thermodynamic equilibrium if for every material point there exists a neighborhood large enough to observe thermodynamic properties, yet small enough for the properties to be essentially non-varying inside the neighborhood, see, e.g., [13]. The size of these neighborhoods corresponds to a macroscopic scale at which the dynamics of the thermodynamic properties can be observed and also modeled.
The first law of thermodynamics restates a fact known from mechanics, namely that energy is a conserved quantity. However, instead of considering directly the mechanical and electromagnetic energy associated to microscopic degrees of freedom, a thermodynamic model with local equilibrium resorts to a phenomenological description. This so-called internal energy is considered to depend on a number of extensive properties, including entropy, and possibly further mechanical or electromagnetic state variables which are observable on the considered macroscopic scale. These quantities then serve as the state variables of a coarse-grained description. Moreover, temperature is defined as the derivative of the internal energy with respect to entropy. Similarly, pressure is given by the derivative with respect to volume and the chemical potential is the derivative with respect to mass, see, e.g., [12].
Models in classical mechanics and electromagnetism exhibit reversible dynamics, meaning that one could in principle arrive again at the initial condition by formally reversing the direction of time (although for many systems this would be practically impossible due to chaos). Incomplete knowledge about the microscopic state immediately leads to an irreversible dynamics and an irreversible dynamics in turn leads to a growth of entropy, as information (e.g. about the initial state of the system) is lost as the system evolves. This is reflected by the second law of thermodynamics which states that the entropy of an isolated system never decreases and that the system evolves towards a state of maximum entropy, see, e.g., [13]. In other words, the uncertainty about the microscopic state can only grow, not decrease.
For models with local equilibrium, the irreversible dynamics can be expressed in terms of so-called thermodynamic fluxes that determine instantaneous changes of coarse-grained state variables. The fluxes are expressed as a function of so-called thermodynamic forces, which are given by local differences in intensive quantities, see, e.g., [13, 14]. For instance, temperature differences cause heat flux, according to Fourier’s law, and differences in chemical potential cause a diffusive mass flux, according to Fick’s law. Further, one may observe cross effects such as a diffusive mass flux caused by temperature differences (Soret effect) and likewise a heat flux caused by differences in chemical potential (Dufour effect). Onsager’s reciprocal relations state that the relation between fluxes and forces must possess a certain symmetry [15]. Specifically, the ratio of the mass flux caused by the Soret effect and the respective temperature difference is equal to the ratio of the heat flux caused by the Dufour effect and the respective difference in chemical potential.
1.5 Related work
Figure 3 gives a high-level view of this subsection.
Figure 3: Informal Venn diagram illustrating that EPHS integrates ideas from graphical and energy-based modeling of physical systems with bond graphs (see, e.g., [7, 8]), the metriplectic/GENERIC framework for nonequilibrium thermodynamics (see, e.g., [5, 13]), port-Hamiltonian theory for open, dissipative systems (PHS) (see, e.g., [4]), and applied category theory (ACT) research on the formalization of graphical languages and compositional dynamical systems (see, e.g., [9, 16, 17, 18, 19]).
1.5.1 Bond graphs
Bond graphs provide a graphical notation for representing multi-domain models, particularly those involving mechanical and electrical systems, see, e.g., [7, 8]. The main idea is that elements interact via so-called bonds that represent links for energy exchange. Each bond is associated with two dynamic variables, called flow and effort, whose product gives the exchanged power. Bonds can be connected through two types of junctions: 00-junctions, where the flow variables of all connected ports balance while all effort variables are equal, and $1$-junctions, where the effort variables balance while all flow variables are equal. In both cases, the net power at a junction is zero. The elements are commonly generalized capacitors, inductors, transformers, resistors, etc. The junctions are then seen as generalized Kirchhoff’s circuit laws.
1.5.2 Port-Hamiltonian systems
Port-Hamiltonian Systems (PHS) were originally inspired by relating the dynamical systems obtained from generalized bond graphs111 While bond graphs commonly use generalized capacitors and inductors as storage elements, so-called generalized bond graphs use only generalized capacitors together with gyrator elements [21]. to Hamiltonian mechanics [20]. While many authors make use of PHS in input-state-output form, PHS generally constitute an acausal systems theory, also inspired by Willems’ behavioral approach [22, 23]. The dynamics of networks composed of energy storing elements and reversible coupling elements (generalized transformers and gyrators) can, from a geometric perspective, be formalized as Hamiltonian systems whose Poisson structure reflects the network topology [24]. The generalization to systems with algebraic constraints and/or external port variables is achieved using Dirac structures [25]. These offer a relational approach that unifies Poisson and presymplectic structures (see, e.g., [1, 26]). By modeling resistive elements as relations imposed on external port variables, Hamiltonian dynamics can be combined with (dissipative) gradient dynamics. A key feature of PHS is their composability: interconnecting simpler systems yields a more complex one. This is achieved by composing the Dirac structures of the individual systems with an ‘interconnecting Dirac structure’ (see, e.g., [4]). Another key feature is that every PHS satisfies a power-balance equation, stating that at every instant the sum of the stored energy and the dissipated energy equals the supplied energy. As the dissipated energy cannot be negative, PHS are inherently passive. In this context, the term ‘energy’ stems from the ‘Hamiltonian’ storage function. For EPHS, this function is interpreted as the stored exergy, making the power-balance equation a statement about rates of exergy transfer. This interpretation fully reconciles the inherent passivity of PHS with the first and second law of thermodynamics, as exergy is indeed dissipated (or destroyed) by irreversible dynamics.
1.5.3 Metriplectic and GENERIC framework
Metriplectic systems as well as the general equation for reversible-irreversible coupling (GENERIC) refer to a framework for combining reversible and irreversible dynamics in a thermodynamically consistent manner, see, e.g. [5, 27, 28, 29, 30]. The reversible part is modeled as a Hamiltonian dynamics, which inherently conserves energy, while the irreversible part is modeled as a (generalized) gradient flow, which inherently drives the system toward maximum entropy. Two additional conditions ensure that the reversible dynamics conserves entropy, and the irreversible dynamics conserves energy. As a result, the combined dynamics respects both the first law and the second law (see, e.g., [13, 6]). While many well-known physical models can be identified as metriplectic systems or instances of the GENERIC, the framework is particularly useful as a guide for developing new thermodynamic models.
1.5.4 Applied category theory
Category theory was first applied in algebraic geometry and topology and now is used to organize many branches of mathematics (see, e.g., [10]). In recent decades, category theory has grown beyond pure math to organize scientific and engineering disciplines, giving birth to applied category theory (see, e.g. [9]). The following ideas from categorical systems theory (see, e.g., [19]) have inspired our work.
  1. 1.
    The shift of focus from individual systems to how systems relate to each other.
  2. 2.
    An isolated system is a system with empty interface.
  3. 3.
    The description of systems involves syntax and semantics. Syntax consists of combinatorial data that can be manipulated computationally, while semantics are given by geometric objects that exist only platonically. An explicit understanding of the translation from syntax to semantics serves as a guide for computer implementation.
A prominent line of research formalizes graphical languages such as block diagrams [17], Petri nets [31], junctions of bond graphs [32], and passive linear circuits [33]. In each case, the graphical syntax is defined as some sort of string diagrams, using a generators and relations approach.
Another line of research uses the theory of operads to define a graphical syntax that supports hierarchical (de-)composition (see, e.g., [34, 35]). Operads organize formal operations with finitely many inputs (subsystems) and one output (composite system) (see, e.g., [36, 37]). As proposed in [38], the EPHS syntax, is based on the operad of undirected wiring diagrams, which initially appeared in [16], together with a relational semantics used to express database queries, see also [39, 40]. Directed wiring diagrams have been used to formalize nested systems of Moore machines and ODEs with inputs and outputs [41, 42]. A framework for discrete and continuous dynamical systems with directed and undirected notions of composition is presented in [43]. The composition of lossless PHS is studied from a categorical perspective in [44].
1.6 Outline
Section 2 defines system interfaces. Based on this, Section 3 defines interconnection patterns. Section 4 shows how an interconnection pattern implies a relation between the port variables of its interfaces. Section 5 discusses how exergy quantifies the energy stored in a system that is available for doing work. Based on this, Section 6 introduces the primitive systems. Section 7 defines composite systems and how they are formed using the graphical syntax. Section 8 shows that EPHS models are thermodynamically consistent. Section 9 presents an example of a direct current (DC) motor model, along with a comparison to a classical bond graph. Section 10 concludes with a discussion.
2 Interfaces
While the narrative of discovery, usually proceeds from systems to interfaces, logically speaking, the syntax needs to be introduced before the semantics and hence we start with interfaces. Graphically, the interface of a system is a box with a set of ports drawn around it. While we could treat such sets abstractly, as is often done in applied category theory, for practical reasons we instead treat them as sets of names called namespaces.
2.1 Namespaces
Definition 2.1.
A string is a non-empty list of characters and a name is a list of strings. ∎
We write names using a monospace font and with strings separated by dots. For instance, the name consisting of the list of strings “oscillator”, “spring”, “q” is written as $\mathtt{oscillator.spring.q}$. Further, we use [][][ ] to denote the empty name, i.e. the empty list of strings. Given two names $n$ and $n^{\prime}$, we denote their concatenation by $n\ast n^{\prime}$. For example, $\mathtt{a.b}\ast\mathtt{c}=\mathtt{a.b.c}$.
Definition 2.2.
Given two names $p$ and $n$, we say that $p$ is a prefix of $n$ and write $p\leq n$ if there exists any name $s$ such that $n=p\ast s$. In case $s\neq[]$, we say that $p$ is a strict prefix of $n$ and write $p<n$. ∎
Example 2.3.
It holds that $\mathtt{a.b}<\mathtt{a.b.c}$ and $\mathtt{a.b}\leq\mathtt{a.b}$. Also, $[]<\mathtt{a}$. It does not hold that $\mathtt{a.a}\leq\mathtt{a}$. ∎
Definition 2.4.
A set of names $N$ is prefix-free if for all $n,n^{\prime}\in N$, $n$ is not a strict prefix of $n^{\prime}$. A namespace is a finite, prefix-free set of names. Let $\mathsf{Namespace}$ refer to the set of all namespaces. ∎
Example 2.5.
The set of names $\{\mathtt{a},\,\mathtt{b}\}$ is a namespace and so is $\{\mathtt{a.a},\,\mathtt{a.b}\}$. Since $\mathtt{a}<\mathtt{a.b}$, the set of names $\{\mathtt{a},\,\mathtt{a.b}\}$ is not prefix-free and hence not a namespace. ∎
Sets of names are equivalent to prefix trees, called tries in computer science [45].
Example 2.6.
The namespace $\{\mathtt{a.b.b},\,\mathtt{a.b.c},\,\mathtt{b.a},\,\mathtt{b.y},\,\mathtt{c}\}$ corresponds to the following trie:
[Uncaptioned image][Uncaptioned $\qed$
Note that for a namespace $N$, the leaf nodes of its trie correspond to the names in $N$, while for a set of names that is not prefix-free, at least one name corresponds to an internal node. In the following, we think of namespaces both as sets of names and as tries.
Given several namespaces, we could take their union as sets. However, this runs the risk of name conflicts, as it is possible for a given name in the union to have come from either of the original sets. It is for this reason that we introduce the concept of a named sum, which can be thought of as a practical implementation of disjoint union.
Definition 2.7.
Given a namespace $N$, and a function $f\colon N\to\mathsf{Namespace}$, the named sum of $f$ is
$\sum f\>=\>\Bigl{\{}\,n\ast a\,\mid\,n\in N,\>a\in f(n)\,\Bigr{\}}\,.\qed$
We also sometimes write expressions like $\sum\left[\mathtt{a}\mapsto I_{1},\,\mathtt{b}\mapsto I_{2}\right]$, which refers to the named sum of the anonymous function defined by
$x\>\mapsto\>\begin{cases}I_{1}&\text{if }x=\mathtt{a}\\ I_{2}&\text{if }x=\mathtt{b}\,.\end{cases}$
Example 2.8.
Given $N=\{\mathtt{p},\,\mathtt{q}\}$, we have
$\sum\left[\mathtt{l}\mapsto N,\,\mathtt{r}\mapsto N\right]\>=\>\bigl{\{}\mathtt{l.p},\,\mathtt{l.q},\,\mathtt{r.p},\,\mathtt{r.q}\bigr{\}}\,.\qed$
Proposition 2.9.
Given a namespace $N$ and a function $f\colon N\to\mathsf{Namespace}$, the named sum $\sum f$ is again a namespace, that is to say, it is prefix-free. ∎
Proof.
The trie $\sum f$ is simply the trie $N$ with, for each leaf node $n\in N$, the trie $f(n)$ grafted onto it. ∎
As shown next, we can use namespaces and named sums to develop a simple and effective formalism for working with hierarchically-structured data.
2.2 Directories
The interface of a system is given by a namespace, where each name refers to a port. However, this is not all, as each port has two associated attributes. This motivates the definition of directories:
Definition 2.10.
Given a set $A$ (of possible attributes), we define
$\mathsf{Dtry}(A)\>=\>\Bigl{\{}\,(N,\,\tau)\,\mid\,N\in\mathsf{Namespace},\,\tau\colon N\to A\,\Bigr{\}}\,.$
An element $(N,\,\tau)\in\mathsf{Dtry}(A)$ is called a directory of elements of $A$. We also just write $\tau\in\mathsf{Dtry}(A)$ with $N=\operatorname{dom}(\tau)$. ∎
The name ‘directory’ points to the similarity with hierarchical file systems, where complete paths (names) are associated to file contents (attributes).
A directory $(N,\,\tau)\in\mathsf{Dtry}(A)$ can be seen as a trie $N$, where each leaf node $n\in N$ has an associated attribute $\tau(n)\in A$. Hence, the concept of named sum carries over from $\mathsf{Namespace}$ to $\mathsf{Dtry}(A)$, for any fixed set $A$. Named sum can then be seen as a flattening operation, analogous to flattening a list of lists. Specifically, for the set of all lists of elements of $A$
$\mathsf{List}(A)\>=\>\Bigl{\{}\,(n,\,f)\,\mid\,n\in\mathbb{N},\,f\colon\{1,\,\ldots,\,n\}\to A\,\Bigr{\}}\,,$
flattening is a map $\mathsf{List}(\mathsf{List}(A))\to\mathsf{List}(A)$, which for instance sends [[1, 2],[3, 4]]1234[[1,\,2],\,[3,\,4]][ [ 1 , 2 ] , [ 3 , 4 ] ] to [1, 2, 3, 4]1234[1,\,2,\,3,\,4][ 1 , 2 , 3 , 4 ]. Similarly, named sum is a map $\mathsf{Dtry}(\mathsf{Dtry}(A))$ to $\mathsf{Dtry}(A)$. An input to this is a trie $N$, which has associated to each leaf node $n\in N$ another trie $N_{n}$, whose leaves carry elements of $A$. The named sum operation then simply grafts onto each node $n\in N$ its associated trie $N_{n}$, carrying along the elements of $A$ at the leaves. In terms of category theory, the flattening operation, together with a unit operation, makes $\mathsf{List}$ and $\mathsf{Dtry}$ examples of monads.
2.3 Bundle of quantities
The first attribute associated to a port represents the physical quantity that can be exchanged via the port, e.g. momentum or entropy, along with a state space in which values of the quantity live. For instance a momentum variable can take values in $\mathbb{R}$ if motion along a single axis is considered, or it can take values in the vector space $\mathfrak{se}^{*}(3)$ if rigid body motions in $3$-dimensional Euclidean space are considered, etc.
Definition 2.11.
Let $\mathbf{S}$ be a finite set of state spaces (i.e. smooth manifolds) and for each space $\mathcal{X}\in\mathbf{S}$, let $\mathbf{Q}_{\mathcal{X}}$ be a finite set of physical quantities taking values in $\mathcal{X}$. Further, let $\mathbf{Q}$ be the disjoint union of all quantities, i.e.
$\mathbf{Q}\>=\>\bigsqcup_{\mathcal{X}\in\mathbf{S}}\mathbf{Q}_{\mathcal{X}}\,.$
This forms a bundle of quantities (in the category of finite sets), whose projection
$\mathrm{space}:\mathbf{Q}\twoheadrightarrow\mathbf{S}$
returns the underlying space of a quantity. ∎
Example 2.12.
The bundle of quantities given by
$\begin{split}\mathbf{S}\>&=\>\{\mathbb{R},\,\mathrm{SE}(3),\,\mathfrak{se}^{*}(3)\}\\ \mathbf{Q}_{\mathbb{R}}\>&=\>\{\mathtt{displacement},\,\mathtt{momentum},\,\mathtt{entropy},\,\mathtt{volume}\}\\ \mathbf{Q}_{\mathrm{SE}(3)}\>&=\>\{\mathtt{pose}\}\\ \mathbf{Q}_{\mathfrak{se}^{*}(3)}\>&=\>\{\mathtt{momentum}\},\end{split}$
would be sufficient for some spatially-lumped models from mechanics and thermodynamics. A port with associated quantity $q=(\mathbb{R},\,\mathtt{entropy})\in\mathbf{Q}$ can exchange entropy and its associated state space is $\mathrm{space}(q)=\mathbb{R}\in\mathbf{S}$. ∎
We note that $(\mathbb{R},\,\mathtt{energy})\notin\mathbf{Q}$, because in the context of EPHS, internal energy is a function of the state variable entropy, and not vice versa, see also [3].
To distinguish reversible and irreversible dynamics later on in Sections 6.3 and 6.4, we make use of the concept of parity, see [46, 3] for details.
Definition 2.13.
The parity of each quantity with respect to time-reversal transformation is given by a function $\mathrm{P}:\mathbf{Q}\rightarrow\{-1,\,+1\}$. ∎
Example 2.14.
For the bundle of quantities from 2.12, we have
$\mathrm{P}(\mathbb{R},\,\mathtt{displacement})=+1$
$\mathrm{P}(\mathbb{R},\,\mathtt{momentum})=-1$
$\mathrm{P}(\mathbb{R},\,\mathtt{entropy})=+1$
$\mathrm{P}(\mathbb{R},\,\mathtt{volume})=+1$
because displacement, entropy, and volume do not change instantaneously when a recording is suddenly played backwards (positive parity), whereas momentum changes its sign (negative parity). ∎
A comprehensive collection of physical quantities with their associated parities would be part of a software implementation.
2.4 Interfaces as directories
Each port has two attributes: the previously discussed quantity, and a Boolean attribute that indicates whether the port is a power port or a state port. Indeed, only power ports allow for an exchange of energy, based on an exchange of the respective quantity. State ports are required whenever the behavior of one system depends on the state of another system, although no energy is exchanged between the two.
Definition 2.15.
An interface is a directory of elements of $\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}$, i.e. let
$\mathsf{Interface}\>=\>\mathsf{Dtry}\bigl{(}\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}\bigr{)}\,.$
Here, element $\mathsf{p}$ stands for power port and $\mathsf{s}$ stands for state port. ∎
Example 2.16.
Let $I=(N,\,\tau)$ be an interface with namespace $N=\{\mathtt{p_{1}},\,\mathtt{p_{2}},\,\mathtt{q}\}$ and attributes given by $\tau(\mathtt{p_{1}})=\tau(\mathtt{p_{2}})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$ and $\tau(\mathtt{q})=\left((\mathbb{R},\,\mathtt{displacement}),\,\mathsf{s}\right)$. Ports $\mathtt{p_{1}}$ and $\mathtt{p_{2}}$ are power ports that can exchange kinetic energy by exchanging momentum, along with information about the current amount of momentum. Port $\mathtt{q}$ is a state port that can only exchange information about the current displacement. ∎
2.5 Bundles of port variables
A state port has one port variable called state. A power port has two additional port variables called flow and effort. Given an interface, the port variables of all its ports combined take values in a vector bundle, called the bundle of port variables associated to the interface.
Definition 2.17.
Let $I=(N,\,\tau)$ be an interface and let $p\in I$ be a port, i.e. $p\in N$.
The state space associated to port $p$ is
$\mathcal{X}_{p}\>=\>(\mathrm{space}\circ\mathrm{pr}_{1}\circ\tau)(p)\,,$
where $\mathrm{pr}_{1}$ projects onto the first attribute, i.e. the quantity.
The state space associated to interface $I$ is given, up to isomorphism, by the Cartesian product
$\mathcal{X}_{I}\>=\>\prod_{p\,\in\,I}\mathcal{X}_{p}\,.$
To define $\mathcal{X}_{I}$ strictly, we assume that the elements of $I$ are in lexicographic order and that the Cartesian product preserves this order. ∎
Example 2.18.
Let $I_{\text{gas}}=(N_{\text{gas}},\,\tau_{\text{gas}})$ be an interface with namespace $N_{\text{gas}}=\{\mathtt{s},\,\mathtt{v}\}$ and attributes given by $\tau_{\text{gas}}(\mathtt{s})=\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$, $\tau_{\text{gas}}(\mathtt{v})=\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$. Then, its associated state space is $\mathcal{X}_{I_{\text{gas}}}=\mathcal{X}_{\mathtt{s}}\times\mathcal{X}_{\mathtt{v}}=\mathbb{R}\times\mathbb{R}$. Let $I_{\text{mass}}=(N_{\text{mass}},\,\tau_{\text{mass}})$ be another interface with $N_{\text{mass}}=\{\mathtt{p}\}$ and $\tau_{\text{mass}}(\mathtt{p})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$. Then, the named sum
$I\>=\>(N,\,\tau)\>=\>\sum[\mathtt{gas}\mapsto I_{\text{gas}},\,\mathtt{mass}\mapsto I_{\text{mass}}]$
has the namespace $N=\{\mathtt{gas.s},\,\mathtt{gas.v},\,\mathtt{mass.p}\}$ with attributes given by $\tau(\mathtt{gas.s})=\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$, $\tau(\mathtt{gas.v})=\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$, $\tau(\mathtt{mass.p}))=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$. We note that
$\mathcal{X}_{I}\>=\>\mathcal{X}_{I_{\text{gas}}}\times\mathcal{X}_{I_{\text{mass}}}\,.\qed$
Given any interface $I$, each port $p\in I$ has an associated state variable $p\mathtt{.x}\in\mathcal{X}_{p}$. If $p$ is a power port, it additionally has a flow variable $p\mathtt{.f}$ and an effort variable $p\mathtt{.e}$. All port variables implicitly depend on time. As detailed in Section 6.1, the flow variable is the instantaneous rate of change of the state variable and the effort variable is the differential of a function of the state variable. Hence, we have $\left(p\mathtt{.x},\,p\mathtt{.f},\,p\mathtt{.e}\right)\in\mathbb{T}\mathcal{X}_{p}$, where $\mathbb{T}\mathcal{X}_{p}=\mathrm{T}\mathcal{X}_{p}\oplus\mathrm{T}^{*}\mathcal{X}_{p}$ denotes the direct sum (or Whitney sum) of the respective tangent and cotangent bundles (see, e.g., [47]). The flow and effort variables are called power variables because their duality pairing $\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle$ yields the instantaneous power that is exchanged through port $p$.
Now we can formally define the bundle of port variables associated to an interface:
Definition 2.19.
Let $I$ be an interface. Based on the Boolean attribute of its ports, we can think of its associated state space $\mathcal{X}_{I}$ as the Cartesian product $\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$, where $\mathcal{X}_{I,p}$ is the state space of all power ports and $\mathcal{X}_{I,s}$ is the state space of all state ports. Accordingly, we let $x=(x_{p},\,x_{s})\in\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$ denote all state variables, and we let $(f,\,e)\in\mathbb{T}_{x_{p}}\mathcal{X}_{I,p}$ denote the flow and effort variables. The bundle of port variables $\mathrm{pr}_{I}:\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ associated to interface $I$ is defined, up to isomorphism, as the pullback bundle of $\mathbb{T}\mathcal{X}_{I,p}$ along the projection $\mathrm{pr}_{p}:\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}\twoheadrightarrow\mathcal{X}_{I,p}$. Hence, $\mathcal{P}_{I}\cong\mathbb{T}\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$.
[…]
Again, to define the bundle of port variables strictly, we assume that the elements of $I$ are in lexicographic order and that the port variables are written in the same order as the ports. ∎
Example 2.20.
Consider again the interface from 2.16. Its bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ has the total space (or port space)
$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\>\ni\>\bigl{(}(\mathtt{p_{1}.x},\,\mathtt{p_{1}.f},\,\mathtt{p_{1}.e}),\,(\mathtt{p_{2}.x},\,\mathtt{p_{2}.f},\,\mathtt{p_{2}.e}),\,\mathtt{q.x}\bigl{)}$
and the base space (or state space)
$\mathcal{X}_{I}\>=\>\mathbb{R}\times\mathbb{R}\times\mathbb{R}\>\ni\>\left(\mathtt{p_{1}.x},\,\mathtt{p_{2}.x},\,\mathtt{q.x}\right).\qed$
3 Syntax
As introduced in Section 1.2, an interconnection pattern expresses how a composite system is formed from a finite number of subsystems. Syntax refers to the entirety of such expressions and their composition, i.e. their hierarchical nesting.
3.1 Interconnection patterns
Figure 4: Interconnection pattern of an isothermal oscillator model. Box $\mathtt{pkc}$ represents the reversible coupling of the potential energy domain represented by the junction on its left and the kinetic energy domain represented by the junction on its right. Box $\mathtt{env}$ represents the isothermal environment, which absorbs the heat that is generated by the damping. The outer box represents the interface of the composite system. Its port $\mathtt{p}$ exposes the kinetic energy domain, allowing for an external forcing of the oscillator model.
Figure 4 shows an interconnection pattern. The mathematical content of such a diagram is captured by the following definition.
Definition 3.1.
An interconnection pattern $(\beta,\,I,\,J)$ is defined by the following data:
  1. 1.
    A directory of interfaces $\beta\in\mathsf{Dtry}(\mathsf{Interface})$. Names in $\operatorname{dom}(\beta)$ refer to inner boxes, $\sum\beta$ is called inner interface and names in $\operatorname{dom}(\sum\beta)$ refer to inner ports.
  2. 2.
    An interface $I$ called outer interface, whose ports are called outer ports.
  3. 3.
    A partition $J$ of the combined interface $P=\sum\left[\mathtt{inner}\mapsto\sum\beta,\,\mathtt{outer}\mapsto I\right]$ such that each subset or junction $j\in J$ contains
    • (a)
      at least one inner port,
    • (b)
      at most one outer port,
    • (c)
      and all ports in $j$ have the same associated quantity. ∎
To prevent name conflicts, the definition distinguishes inner ports and outer ports with the prefixes $\mathtt{inner}$ and $\mathtt{outer}$. While preventing name conflicts is crucial for implementation of the modeling language, we omit these prefixes for conciseness in what follows.
We graphically depict interconnection patterns as follows. Each element $b\in\operatorname{dom}(\beta)$ is the name of an inner box, which we draw as a circle. The ports of its interface $\beta(b)$ are drawn as lines emanating from the box. The ports of the outer interface $I$ are drawn as lines to the surrounding outer box. Each subset $j\in J$ corresponds to a junction drawn as a black dot. The elements of $j\subseteq P$ are precisely the connected ports. To avoid confusion, we recall that the identifier isothermal_osc written on the top left of the outer box refers to the entire pattern (akin to a function name in programming), rather than just to the outer box (which similar to the return value of a function does not need a name).
The pattern from Figure 4 is redrawn in Figure 5, firstly to emphasize that the color of the inner boxes is merely an annotation hinting at the nature of the considered subsystems and secondly to introduce an abbreviated way of writing port names.
Figure 5: Interconnection pattern of an isothermal oscillator model. Whenever all ports connected to a certain junction have the same name, we write the name only once at the junction.
Example 3.2.
For the pattern $(\beta,\,I,\,J)$ depicted in Figure 5, the namespace of inner boxes is
$\operatorname{dom}(\beta)\>=\>\{\mathtt{damping},\,\mathtt{env},\,\mathtt{mass},\,\mathtt{pkc},\,\mathtt{spring}\}\,.$
The function $\beta:\operatorname{dom}(\beta)\rightarrow\mathsf{Interface}$ is given by
$\begin{split}\mathtt{damping}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{damping}})\\ \mathtt{env}\>&\mapsto\>(\{\mathtt{s}\},\,\tau_{\text{env}})\\ \mathtt{mass}\>&\mapsto\>(\{\mathtt{p}\},\,\tau_{\text{mass}})\\ \mathtt{pkc}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{q}\},\,\tau_{\text{pkc}})\\ \mathtt{spring}\>&\mapsto\>(\{\mathtt{q}\},\,\tau_{\text{spring}})\end{split}$
with $\tau_{\text{damping}}(\mathtt{p})=((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})$, etc. Hence, the namespace of the inner interface $\sum\beta$ is
$\operatorname{dom}(\textstyle\sum\beta)\>=\>\{\mathtt{damping.p},\,\mathtt{damping.s},\,\mathtt{env.s},\,\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{pkc.q},\,\mathtt{spring.q}\}\,.$
The namespace of the outer interface $I$ is $\{\mathtt{p}\}$. The partition $J$ is given by
$\bigl{\{}\{\mathtt{damping.p},\,\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{p}\},\,\{\mathtt{pkc.q},\,\mathtt{spring.q}\},\,\{\mathtt{damping.s},\,\mathtt{env.s}\}\bigr{\}}\,.\qed$
We may also denote an interconnection pattern using the condensed notation
${\sum\left[\mathtt{I_{1}}\mapsto I_{1},\,\ldots,\,\mathtt{I_{n}}\mapsto I_{n}\right]}$${J}$${I.}$
In this we implicitly let $\operatorname{dom}(\beta)=\{\mathtt{I_{1}},\,\ldots,\,\mathtt{I_{n}}\}$ and $\beta(\mathtt{I_{i}})=I_{i}$. The left arrow is a surjection because each junction has at least one connected inner port and the right arrow is an injection because each junction has at most one connected outer port. Interconnection patterns are a special case of undirected wiring diagrams [16], which do not have the injectivity and surjectivity conditions.
Multiports are just multiple ports with a common prefix. As demonstrated in Figure 2, this enables a syntactic sugar for expressing parallel connections in a tidy and efficient manner. We defer the details to later work, as the examples considered here make no use of this.
3.2 Composition of interconnection patterns
Whenever the outer interface of one pattern and the interface corresponding to an inner box of another pattern are equal (up to a given renaming of the ports), the former pattern can be substituted into the latter, with no further data required. This is called composition. Having a compositional syntax facilitates dealing with complex systems, as it enables the encapsulation of subsystems. These can be easily reused and replaced, as long as interfaces match. As an example, the isothermal oscillator model can be refined into a nonisothermal oscillator model based on the damper model in Figure 6.
Figure 6: Interconnection pattern of a nonisothermal damper model. Box $\mathtt{mf}$ represents a mechanical friction model connected to the kinetic energy domain represented by the outer port $\mathtt{p}$. Box $\mathtt{tc}$ represents a thermal capacity, which stores the heat generated by friction. Box $\mathtt{ht}$ represents a heat transfer model, which allows the thermal capacity to reach equilibrium with the thermal energy domain represented by the outer port $\mathtt{s}$.
Figure 7: Interconnection pattern of a nonisothermal oscillator model. Except for the graphical layout, the pattern is uniquely defined as the composite of the patterns in Figures 5 and 6.
Substituting the pattern in Figure 6 into the pattern in Figure 5, yields the pattern in Figure 7. Graphically, the former pattern is inserted into the inner box called $\mathtt{damping}$ of the latter pattern. The common interface is deleted and for each of its ports, the two assigned junctions (one in each pattern) are identified. The named sum for interfaces is used to obtain the inner interface of the composed pattern.
Category theory provides a language to talk about composition, and it includes well-behavedness conditions, which ensure that composition works in the way we expect, just as we need not worry about the difference between $(1+2)+3$ and $1+(2+3)$. A future paper shall formalize the composition of interconnection patterns based on directories and multicategories/operads. To limit the scope of this article, we just mention that it is not difficult to see why the procedure for composition outlined in the previous paragraph preserves the three conditions in 3.1.
4 Semantics of interconnection patterns
In Section 2.5 we have defined bundles of port variables associated to interfaces. In Section 3.1 we have defined interconnection patterns, which have an inner interface (containing all ports of the subsystems) and an outer interface (containing the ports of the composite system). We now discuss how to interpret an interconnection pattern as imposing a relation between the port variables of its inner interface and the port variables of its outer interface. We do this in two parts. In the first part, we talk about the relation as a collection of equations, and we discuss its physical interpretation. In the second part, we reify it as a geometric object, namely as a relation between bundles of port variables.
4.1 Interconnection patterns as equations
The semantics of an interconnection pattern $(\beta,\,I,\,J)$ can be seen as a collection of equations. Specifically, for each junction $j\in J$, we have the following.
Let $j_{p}\subseteq j$ be the subset of connected power ports.
  • 1.
    Equality of state. For all ports $p,p^{\prime}\in j$,
    $p\mathtt{.x}\>=\>p^{\prime}\mathtt{.x}$
  • 2.
    Equality of effort. For all power ports $p,p^{\prime}\in j_{p}$,
    $p\mathtt{.e}\>=\>p^{\prime}\mathtt{.e}$
  • 3.
    Equality of net flow.
    $\sum_{p\,\in\,j_{p}\cap\sum\beta}p\mathtt{.f}\>=\>\sum_{p\,\in\,j_{p}\cap I}p\mathtt{.f}$
We note that the sum on the left-hand side has at least one term (surjectivity condition) and the sum on the right-hand side has at most one term (injectivity condition).
As detailed in Section 7, combining these equations with equations that determine the semantics of the subsystems filling the inner boxes yields a collection of equations that determines the semantics of the resulting composite system.
Since all ports connected to a junction have the same quantity, their associated state variables take values in the same space, making equality of state a well-typed equation. As a consequence of this, all effort variables take values in the same fiber of the same cotangent bundle, making equality of effort also well-typed. Similarly, all flow variables take values in the same fiber of the same tangent bundle, making equality of net flow well-typed.
Equality of state is required, since the behavior of one system may in general depend on the state of another system, even if the two systems do not exchange energy. We refer to Figures 8 and 9 for an example of this.
As a consequence of equality of effort and equality of net flow, at every junction $j\in J$ the following power balance equation is satisfied:
$\sum_{p\,\in\,j_{p}\cap\sum\beta}\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle\>=\>\sum_{p\,\in\,j_{p}\cap I}\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle$
Summing over all junctions on both sides gives that the net power supplied to all subsystems is equal to the net power supplied to the composite system.
Junctions correspond to energy domains and so do (outer) ports, since they expose energy domains. Whenever two ports are connected to the same junction the respective energy domains are identified. This is exactly how interconnection patterns compose. Interconnection means that systems share energy domains. The state of a shared energy domain is given by the shared state variable, according to the equality of state. The connected systems may exchange the respective physical quantity according to the equality of net flow. Moreover, based on the equality of effort, this exchange implies a transfer of (exergetic) power.
Example 4.1.
Let’s again consider the pattern in Figure 7. At the junction which represents the kinetic energy domain, the momentum variable (state) is shared and forces (flows) balance at equal velocity (effort):
[…]
Table 1 summarizes the physical interpretation of junctions representing a number of different energy domains.
energy domain state variable flows balance at equal efforts
kinetic momentum forces balance at equal velocity
magnetic flux linkage voltages balance at equal current
potential displacement velocities balance at equal force
hydraulic volume volume rates balance at equal pressure
electric charge currents balance at equal voltage
thermal entropy entropy rates balance at equal temperature
Table 1: Physical interpretation of junctions as energy domains.
We can now make sense of the surjectivity and injectivity conditions distinguishing interconnection patterns form undirected wiring diagrams. The surjectivity condition says that each junction must have at least one connected inner port. In other words, an energy domain can only belong to a composite system if it belongs to at least one subsystem. The injectivity condition says that each junction may have at most one connected outer port. In other words, an energy domain can be exposed at most once. Exposing an energy domain twice and identifying it again at a higher level by connecting the two respective ports would lead to undefined flow variables.
4.2 Interconnection patterns as relations
[…]
Analogous to the prototypical example of a relation between two sets, the relation assigned to an interconnection pattern is defined by a vector subbundle of the Cartesian product of the bundle of port variables associated to the inner interface and the bundle of port variables associated to the outer interface.
Definition 4.2.
Let $P=(\beta,\,I,\,J)$ be an interconnection pattern and let $I_{\text{in}}=\sum\beta$ denote its inner interface. Further, let $\mathcal{P}_{I_{\text{in}}}\twoheadrightarrow\mathcal{X}_{I_{\text{in}}}$ be the bundle of port variables associated to the inner interface, and let $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ be the bundle of port variables associated to the outer interface. Based on the partition $J$, let $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ be the subspace where equality of state, equality of effort, and equality of net flow are satisfied. Similarly, let $\mathcal{X}_{P}\subset\mathcal{X}_{I_{i}}\times\mathcal{X}_{I}$ be the subspace where equality of state is satisfied. Then, the semantics assigned to interconnection pattern $P$ is the relation $\mathcal{P}_{P}:\mathcal{P}_{I_{\text{in}}}\nrightarrow\mathcal{P}_{I}$ between the port variables of the inner interface and the port variables of the outer interface that is defined by the vector subbundle $\mathcal{P}_{P}\twoheadrightarrow\mathcal{X}_{P}$. ∎
Figure 8: Interconnection pattern of the DC shunt motor model discussed in Section 9. The stator and rotor are connected to a common electric energy domain that is exposed via the outer port $\mathtt{q}$. The kinetic energy domain of the rotor is exposed via the outer port $\mathtt{p}$. The conversion between electric and kinetic energy in the rotor depends on the magnetic energy domain of the stator, which is represented by the junction with two connected state ports named $\mathtt{b_{s}}$.
Example 4.3.
Consider the interconnection pattern $P=(\beta,\,I,\,J)$ depicted in Figure 8 with names of inner boxes given by $\operatorname{dom}(\beta)=\{\mathtt{rotor},\,\mathtt{stator}\}$ and interfaces given by
$\beta(\mathtt{rotor})\>=\>\left(\{\mathtt{b_{s}},\,\mathtt{p},\,\mathtt{q}\},\,\tau_{\text{r}}\right)$
$\beta(\mathtt{stator})\>=\>\left(\{\mathtt{b_{s}},\,\mathtt{q}\},\,\tau_{\text{s}}\right)$
$I\>=\>\left(\{\mathtt{p},\,\mathtt{q}\},\,\tau\right)\,,$
where
$\tau_{\text{r}}(\mathtt{b_{s}})\>=\>\tau_{\text{s}}(\mathtt{b_{s}})\>=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
$\tau_{\text{r}}(\mathtt{p})\>=\>\tau(\mathtt{p})\>=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
$\tau_{\text{r}}(\mathtt{q})\>=\>\tau_{\text{s}}(\mathtt{q})\>=\>\tau(\mathtt{q})\>=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})\,.$
Further, the interconnection is determined by
$J\>=\>\bigl{\{}\{\mathtt{rotor.b_{s}},\,\mathtt{stator.b_{s}}\},\,\{\mathtt{rotor.p},\,\mathtt{p}\},\,\{\mathtt{rotor.q},\,\mathtt{stator.q},\,\mathtt{q}\}\bigr{\}}\,.$
The port space of the inner interface $I_{\text{in}}=\sum\beta$ is
$\begin{split}\mathcal{P}_{I_{\text{in}}}\>=\>&\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\times\mathbb{T}\mathbb{R}\\ \>\ni\>&\bigl{(}\mathtt{rotor.b_{s}.x},\,(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e}),\,\\ &\>\>(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e}),\,\\ &\>\>\mathtt{stator.b_{s}.x},\,(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\bigr{)}\end{split}$
and the port space of the outer interface $I$ is
$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\>\ni\>\bigl{(}(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e}),\,(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\bigr{)}\,.$
Finally, the subspace $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ defining the relation is
$\begin{split}\mathcal{P}_{P}\>=\>\bigl{\{}&\bigl{(}\mathtt{rotor.b_{s}.x}\,,\,(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e})\,,\,\\ &(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e})\,,\,\\ &\mathtt{stator.b_{s}.x}\,,\,(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\,,\,\\ &(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\,,\,(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\bigr{)}\in\mathcal{P}_{I_{i}}\times\mathcal{P}_{I}\,\mid\,\\ &\mathtt{rotor.b_{s}.x}=\mathtt{stator.b_{s}.x}\,,\,\\ &\mathtt{rotor.p.x}=\mathtt{p.x}\,,\,\\ &\mathtt{rotor.p.f}=\mathtt{p.f}\,,\,\\ &\mathtt{rotor.p.e}=\mathtt{p.e}\,,\,\\ &\mathtt{rotor.q.x}=\mathtt{stator.q.x}=\mathtt{q.x}\,,\,\\ &\mathtt{rotor.q.f}+\mathtt{stator.q.f}=\mathtt{q.f}\,,\,\\ &\mathtt{rotor.q.e}=\mathtt{stator.q.e}=\mathtt{q.e}\,\bigr{\}}\,.\qed\end{split}$
Given a hierarchy of interconnection patterns, we can first compose the patterns and then determine the relation assigned to the composed pattern. Alternatively, we can first determine the relation assigned to each individual pattern and then compose these relations. Both ways lead to the same result. The combinatorial syntax and its relational semantics hence compose in compatible ways. In the language of category theory, we say that the (assignment of) semantics is functorial. A future paper shall use directories and the theory of multicategories/operads to demonstrate that 4.2 indeed defines a suitable notion of functor.
5 Exergy
Every system is ultimately determined by a power-preserving interconnection of primitive subsystems, called components. These represent elementary physical behaviors, namely storage, and reversible as well as irreversible exchange of energy. To achieve a thermodynamically consistent combination of reversible and irreversible dynamics, components are defined with respect to an exergy reference environment. Before components can be defined in the next section, in this section we define the reference environment. We start with its physical interpretation.
5.1 Intuition
As briefly discussed in Section 1.4, thermodynamic models have an irreversible dynamics, since not all physically-relevant degrees of freedom are determined by the state variables. According to the first and second law of thermodynamics, internal energy, which is given by an ‘imprecise’ phenomenological description, cannot be fully converted into mechanical or electromagnetic energy, which is ‘precise’, in the sense that all relevant degrees of freedom are fully determined by the state variables.
The irreversible degradation of ‘mechanically precise’ forms of energy into ‘thermodynamically imprecise’ forms of energy can be quantified using the concept of exergy. A hypothetical conversion device, whose mode of operation is only restricted by the first and second law of thermodynamics is used to determine how much mechanical (or electromagnetic) energy can be extracted from a system, before it reaches thermodynamic equilibrium with a reference environment, see Figure 9. The environment is defined by intensive variables, such as temperature and pressure, which are constant in time and space.
Figure 9: To assess the exergy content of a system, it is combined with a reference environment. The exergy content is then defined as the maximum amount of work that can be extracted from the combined system via a hypothetical conversion device that operates completely reversibly.
The ideal conversion device is not restricted to interacting with the system through its interface. Instead, the device operates in a counterfactual where all energy domains of both the system and the environment can be accessed. The exergy content of a composite system is hence obtained by (recursively) summing the exergy stored in each of its storage components. The exergy content of a component storing purely mechanical or electromagnetic forms of energy is simply equal to its energy content. For components storing internal energy, the exergy content is assessed based on the ideal conversion device and the reference environment. A concrete statement of this follows in 6.3, after the definition of the exergy storage function in 6.2.
For more information about exergy, its history, and its links to port-Hamiltonian as well as metriplectic or GENERIC systems, we refer to [3].
5.2 Reference environment
The internal energy of a thermodynamic system is a function of its entropy, and possibly further extensive quantities such as volume, mass of chemical species, etc. To assess the exergy content of the system, we need for each such variable an energy-conjugate, intensive variable attributed to the reference environment. Since these intensive variables are constant, they determine the equilibrium state of the combined system, see Figure 9.
Definition 5.1.
A reference environment $(I_{R},\,\lambda_{R})$ is defined by an interface $I_{R}=(N_{R},\,\tau_{R})$ whose ports are all power ports and have a unique extensive quantity in $\mathbf{Q}_{\mathbb{R}}$. One port must have the quantity $(\mathbb{R},\,\mathtt{entropy})$. Moreover, each port $p\in I_{R}$ has an associated value $\lambda_{R}(p)\in\mathbb{R}^{*}$ of the energy-conjugate intensive quantity. ∎
Uniqueness of quantities, i.e. injectivity of $\tau_{R}:N_{R}\rightarrow\mathbf{Q}_{\mathbb{R}}\times\{\mathsf{p}\}$, is required to determine the exergy storage function of a system from its energy storage function. Regarding the physical interpretation, each port corresponds to an energy domain that is present in the environment. Allowing multiple energy domains of the same kind would conflict with the environment always being in equilibrium with itself.
Example 5.2.
The minimal definition of a reference environment represents an isothermal heat bath. This amounts to an interface with a single entropy port and an associated reference temperature, henceforth denoted by ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$:
$N_{R}$ =$=\>\{\mathtt{s}\}$
$\tau_{R}(\mathtt{s})$ $=\>\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$
$\lambda_{R}(\mathtt{s})$ $=\>{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,.$
To determine the exergy content of pressurized fluids with respect to an isobaric atmosphere, the definition of the environment is extended to $N_{R}=\{\mathtt{s},\,\mathtt{v}\}$ with
$\tau_{R}(\mathtt{v})$ $=\>\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$
$\lambda_{R}(\mathtt{v})$ $=\>{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\,.$
The minus sign in front of the reference pressure ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}}$ stems from the fact that more volume means less energy. To model for instance a mass diffusion process, a port $\mathtt{m}$ representing the mass of a chemical species with a constant chemical potential ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mu_{0}}$ can be included, i.e. let $N_{R}=\{\mathtt{s},\,\mathtt{v},\,\mathtt{m}\}$ with
[…]
A software implementation should already include a comprehensive definition of a reference environment. Allowing the user to change the values of the intensive quantities would be of interest for exergy analysis and thermodynamic optimization.
6 Primitive systems
At the lowest level, any hierarchically-structured model is defined by primitive systems, called components. There are four kinds of them:
  • 1.
    Storage components represent energy storage.
  • 2.
    Reversible components represent reversible energy exchange
    or a transformation between different forms of describing an energy domain.
  • 3.
    Irreversible components represent irreversible energy exchange.
  • 4.
    Environment components may be used to model energy exchange with
    the reference environment.
Components are in general defined with respect to a reference environment $(I_{R},\,\lambda_{R})$. Further, every component has an interface $I=(N,\,\tau)$ with port variables $(x,\,f,\,e)\in\mathcal{P}_{I}$. To define the exergy storage function of storage components and to state conditions that reversible and irreversible components have to satisfy, we define the following notation: For any port $p\in I$, let
$\lambda(p)\>=\>\begin{cases}\lambda_{R}(p_{R})&\exists\,p_{R}\in I_{R}\,\colon\,\tau_{R}(p_{R})\>=\>\tau(p)\\ 0\in\mathrm{T}^{*}\mathcal{X}_{p}&\text{else.}\end{cases}$
With regard to the existential quantifier, this is well-defined due to the injectivity of $\tau_{R}$. For instance, if $p$ represents a thermal energy domain then $\lambda(p)={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$, whereas if $p$ represents a kinetic energy domain then $\lambda(p)=0$, since the environment does not have a kinetic energy domain. We extend the above notation to an entire interface: Given for example $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
$\lambda(I)\>=\>\bigl{(}\lambda(\mathtt{p_{1}}),\,\ldots,\,\lambda(\mathtt{p_{n}})\bigr{)}\,.$
We also need a similar notation that references a specific environment port: For any port $p\in I$ and $p_{R}\in I_{R}$, let
$\lambda(p,\,p_{R})\>=\>\begin{cases}\lambda_{R}(p_{R})&\tau_{R}(p_{R})\>=\>\tau(p)\\ 0\in\mathrm{T}^{*}\mathcal{X}_{p}&\text{else.}\end{cases}$
Again, given for example $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
$\lambda(I,\,p_{R})\>=\>\bigl{(}\lambda(\mathtt{p_{1}},\,p_{R}),\,\ldots,\,\lambda(\mathtt{p_{n}},\,p_{R})\bigr{)}\,.$
6.1 Storage components
Storage components are primitive systems that represent energy storage. Examples include elastic springs, movable masses, capacitors, inductors, and thermal capacities.
Definition 6.1.
A storage component $(I,\,E)$ is defined by an interface $I$ containing only power ports and an energy storage function $E:\mathcal{X}_{I}\rightarrow\mathbb{R}$. ∎
The function $E$ yields the stored energy for any state $x\in\mathcal{X}_{I}$. Based on the reference environment $I_{R}$, the exergy content is determined as follows.
Definition 6.2.
Let $(I,\,E)$ be a storage component. Its exergy storage function $H:\mathcal{X}_{I}\rightarrow\mathbb{R}$ is defined by
$H(x)\>=\>E(x)-\sum_{p\,\in\,I}\sum_{\begin{subarray}{c}p_{R}\,\in\,I_{R}\\ \tau_{R}(p_{R})\,=\,\tau(p)\end{subarray}}\langle\lambda_{R}(p_{R})\mid p\mathtt{.x}\rangle\,.\qed$
The exergy function $H$ follows directly from the energy function $E$ by subtracting the pairing between each state variable representing an extensive quantity also present in the environment, and the constant, energy-conjugate quantity from the environment. We come back to this in 6.3.
[…]
The effort variables are given by the differential of the exergy function evaluated at the current state $x\in\mathcal{X}_{I}$, i.e.
$\forall\,p\in I\>:\>p\mathtt{.e}\>=\>\left.\partialderivative{H}{(p\mathtt{.x})}\right|_{x}\,.$
The evolution of the state variables is obtained by integrating the flow variables, i.e.
[…]
Example 6.3.
Consider a storage component $(I_{\text{gas}},\,E_{\text{gas}})$ representing a gas-filled compartment whose entropy and volume can change. Its interface $I_{\text{gas}}$ is defined in 2.18, and we assume knowledge of its energy storage function $E_{\text{gas}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. With $x=(s,\,v)$, the exergy storage function is then given by
$H_{\text{gas}}(s,\,v)\>=\>E_{\text{gas}}(s,\,v)\,-\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,s\,+\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}}\,v\,.$
As the temperature is defined by $\theta=\partialderivative{E_{\text{gas}}}{s}$ and the (negative) pressure is defined by $-\pi=\partialderivative{E_{\text{gas}}}{v}$, the effort variables are given by
$\begin{split}\mathtt{s.e}\>&=\>\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\\ \mathtt{v.e}\>&=\>-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}})\,.\end{split}$
Since the entropy rate $\dot{s}=\mathtt{s.f}$ is related to the heat flow rate $\Phi$ via $\dot{s}=\frac{1}{\theta}\,\Phi$, the exergetic power at the entropy port is $\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle=\frac{\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}{\theta}\,\Phi=\eta_{\text{Carnot}}\,\Phi$, where $\eta_{\text{Carnot}}$ is the efficiency of an ideal heat engine operating between a hot source with temperature $\theta$ and a cold sink with temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$. Considering a piston with the pressurized gas on one side and the atmosphere on the other, $\langle\mathtt{v.e}\mid\mathtt{v.f}\rangle=-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}})\,\dot{v}$ is the power that can be extracted from the pressurized gas when the volume expansion rate is $\dot{v}$. ∎
Next, we formally state the semantics of a storage component with interface $I$ in terms of a relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$. This is used in Section 7, where we define the semantics of a composite system in terms of the semantics of its primitive subsystems and the semantics of its (hierarchy of nested) interconnection pattern(s).
Definition 6.4.
The semantics assigned to storage component $(I,\,E)$ is the relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
[…]
where $H$ is induced from $E$ according to 6.2. ∎
6.2 Environment components
From a user perspective, environment components provide a simple alternative to storage components, when dealing with low-grade energy.They may represent an isothermal heat bath, which absorbs heat generated by irreversible components, or an isobaric atmosphere, which exchanges pressure-volume work with other components, etc. Environment components are essentially storage components with a trivial energy and exergy function.
Definition 6.5.
An environment component is defined by a subinterface $I\subseteq I_{R}$ containing the exposed energy domains of the reference environment. ∎
The state variables keep track of the exchanged entropy, volume, etc. As for storage components, their rate of change is given by the flow variables, i.e.
[…]
By definition, the environment has zero exergy content. Since the effort variables are given by the differential of the exergy function, they are also zero:
[…]
In agreement with the definition, the chain rule also tells us that the rate of change of the stored exergy is zero.
Definition 6.6.
The semantics assigned to environment component $I\subseteq I_{R}$ is the relation
$\mathcal{E}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
[…]
6.3 Reversible components
In a generalized sense, reversible components represent gyrators, transformers, and/or kinematic constraints. Generalized gyrators may describe a coupling of kinetic and potential energy domains in mechanics, or a coupling of magnetic and electric energy domains in electromagnetism. Examples where one domain reversibly interacts with itself are gyroscopic forces in rigid body dynamics and self-advection of kinetic energy in fluid dynamics. The Lorentz force is an example where a kinetic energy domain interacts with itself in a way that depends on the state of a magnetic energy domain. Generalized transformers may for instance describe ideal mechanical levers or electrical transformers. Frame transformations in multibody systems are an example relating different representations of the same energy domain. Examples of kinematic constraints include joint constraints in multibody dynamics and the incompressibility constraint in fluid dynamics. Further, the parallel connection of capacitors, the series connection of inductors or springs, and rigid links between point masses are described by constraints.
The geometric structure behind reversible components is called Dirac structure. For our purposes, a Dirac structure $\mathcal{D}$ is a power-preserving, locally linear relation of port variables. More precisely, we could define $\mathcal{D}\subset\mathcal{P}_{I}$ as a maximally isotropic vector subbundle of the bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$, see [1, 4, 26]. However, to state additional conditions that reversible components have to satisfy, we give a more practical definition of Dirac structures in terms of a particular representation, which might be called constrained hybrid input-output representation (cf. [4]).
Definition 6.7.
Let $I$ be an interface and let $\mathcal{X}_{I,p}\cong\mathcal{X}_{1}\times\mathcal{X}_{2}\ni x_{p}=(x_{1},\,x_{2})$ be a factorization of the state space of all power ports with flow and effort variables correspondingly denoted by $f=(f_{1},\,f_{2})$ and $e=(e_{1},\,e_{2})$. A Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ admits the following representation:
$\mathcal{D}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,\exists\,\lambda_{c}\in V^{*}\,\colon\left[\begin{array}[]{c}f_{1}\\ \hline\cr e_{2}\\ \hline\cr 0\end{array}\right]\,=\,\left[\begin{array}[]{c | c | c}L(x)&-g(x)&C^{*}(x)\\ \hline\cr g^{*}(x)&0&0\\ \hline\cr-C(x)&0&0\end{array}\right]\,\left[\begin{array}[]{c}e_{1}\\ \hline\cr f_{2}\\ \hline\cr\lambda_{c}\end{array}\right]\right\}$
Here, $L:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}\mathcal{X}_{1}$, $g:\mathrm{T}\mathcal{X}_{2}\rightarrow\mathrm{T}\mathcal{X}_{1}$, its linear dual $g^{*}:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}^{*}\mathcal{X}_{2}$,
$C:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathcal{X}_{I}\times V$, and its dual $C^{*}:\mathcal{X}_{I}\times V^{*}\rightarrow\mathrm{T}\mathcal{X}_{1}$ are vector bundle maps over the identity on $\mathcal{X}_{I}$. Further, $V$ is a vector space and $L(x)$ is skew-symmetric for all $x\in\mathcal{X}_{I}$. ∎
The maps $L$, $g$ and $C$ respectively define a generalized gyrator, a generalized transformer and a constraint. The constraint is enforced via the Lagrange multiplier $\lambda_{c}$. Due to the skew-symmetry of the matrix in 6.7, exergetic power is conserved, i.e.
[…]
We remark that this actually defines an almost Dirac structure, as no integrability condition, akin to the Jacobi identity for Poisson structures, is required. However, in the unconstrained or holonomically constrained case, such a condition can be checked.
Example 6.8.
Let the interface $I_{\text{hkc}}=(N_{\text{hkc}},\,\tau_{\text{hkc}})$ be given by
$\begin{split}N_{\text{hkc}}\>&=\>\{\mathtt{v_{1}},\,\mathtt{v_{2}},\,\mathtt{p}\}\\ \tau_{\text{hkc}}(\mathtt{v_{1}})\>&=\>\tau_{\text{hkc}}(\mathtt{v_{2}})\>=\>((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p})\\ \tau_{\text{hkc}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\end{split}$
and let $a\in\mathbb{R}$ be a constant surface area of a piston. Then, the equation
$\left[\begin{array}[]{c}\mathtt{v_{1}.f}\\ \mathtt{v_{2}.f}\\ \mathtt{p.f}\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&-a\\ 0&0&+a\\ +a&-a&0\end{array}\right]\,\left[\begin{array}[]{c}\mathtt{v_{1}.e}\\ \mathtt{v_{2}.e}\\ \mathtt{p.e}\end{array}\right]$
defines a Dirac structure $\mathcal{D}_{\text{hkc}}\subset\mathcal{P}_{I_{\text{hkc}}}$. This describes a generalized gyrator that couples the two hydraulic energy domains at the front and the back side of a piston with the kinetic energy domain of the piston’s moving mass. ∎
A reversible component is defined by a Dirac structure that satisfies two additional conditions ensuring thermodynamic consistency.
Definition 6.9.
A reversible component $(I,\,\mathcal{D})$ is given by an interface $I$ and a Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ such that all quantities present in the reference environment (in particular entropy) are conserved and the time-reversal invariance property holds. The first condition requires that $\forall\,p_{R}\in I_{R}$ and $\forall\,x\in\mathcal{X}_{I}\,\colon$
$0\>=\>L(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>g^{*}(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>C(x)\,\lambda(I_{1},\,p_{R})\,.$
The time-reversal invariance property holds if $\forall\,x\in\mathcal{X}_{I}\,\colon$
${(L(x))}^{ij}\neq 0$ $\implies$-⟹ - $\mathrm{P}(x_{1}^{i})\>$ $=\>\mathrm{P}({(L(x))}^{ij})\,\mathrm{P}(x_{1}^{j})$
${(g(x))}^{ik}\neq 0$$\implies$ $\mathrm{P}(x_{1}^{i})\>$ $=\>\mathrm{P}({(g(x))}^{ik})\,\mathrm{P}(x_{2}^{k})\,.$
Here, we use abstract index notation without implied summation. Indices $i$ and $j$ range over the power ports associated with $\mathcal{X}_{1}$ and $k$ ranges over the power ports associated with $\mathcal{X}_{2}$. The state variable is accordingly split as $x=(x_{p},\,x_{s})=((x_{1},\,x_{2}),\,x_{s})$ and $\mathrm{P}$ yields the parity of the respective quantity with $\mathrm{P}(\mathrm{const.})=+1$. ∎
According to the chain rule, a Dirac structure $\mathcal{D}$ conserves the quantity corresponding to some port $p_{R}\in I_{R}$ if any resulting flow variable $f_{1}$ satisfies $0=\langle\lambda(I_{1},\,p_{R})\mid f_{1}\rangle$. Inserting the expression for $f_{1}$, this condition is satisfied if $0=\langle\lambda(I_{1},\,p_{R})\mid L(x)\,e_{1}\rangle$ is satisfied for any effort variable $e_{1}$, and $0=\langle\lambda(I_{1},\,p_{R})\mid g(x)\,f_{2}\rangle$ is satisfied for any flow variable $f_{2}$, and $0=\langle\lambda(I_{1},\,p_{R})\mid C^{*}(x)\,\lambda_{c}\rangle$ is satisfied for any Lagrange multiplier $\lambda_{c}$. Taking the linear adjoint (transpose), the first condition requires that $0=\langle L^{*}(x)\,\lambda(I_{1},\,p_{R})\mid e_{1}\rangle$ for any $e_{1}$. This simplifies to the first part of the condition in the definition, since $L$ is skew-adjoint. The other two parts follow analogously. Regarding details on the time-reversal invariance condition, we refer to [46] or [3].
Example 6.10.
The reversible component $(I_{\text{hkc}},\,\mathcal{D}_{\text{hkc}})$, as already defined in 6.8, describes the coupling between the kinetic energy domain of a piston with surface area $a$ and the two hydraulic energy domains on either side of the piston. The first condition here ensures that volume is conserved. It is satisfied since
$\left[\begin{array}[]{c}0\\ 0\\ 0\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&-a\\ 0&0&+a\\ +a&-a&0\end{array}\right]\,\left[\begin{array}[]{c}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\\ 0\end{array}\right]\,.$
The time-reversal invariance property holds since
$\begin{split}-\mathrm{P}(\mathtt{(\mathbb{R},\,volume}))\>&=\>\mathrm{P}(-a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ -\mathrm{P}((\mathbb{R},\,\mathtt{volume}))\>&=\>\mathrm{P}(+a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\,.\end{split}$
Checks for the other two non-zero entries of $L$ pass due to skew-symmetry of $L$. ∎
The semantics of a reversible component is simply given by its Dirac structure interpreted as a relation.
Definition 6.11.
The semantics assigned to reversible component $(I,\,\mathcal{D})$ is the relation $\mathcal{D}:1\nrightarrow\mathcal{P}_{I}$ that is given by its Dirac structure $\mathcal{D}$. ∎
The symbol $1$ in $1\nrightarrow\mathcal{P}_{I}$ refers to the Cartesian unit for vector bundles, i.e. the vector bundle with a single point as its base space and the zero-dimensional vector space above it.
6.4 Irreversible components
Irreversible components are phenomenological models of irreversible processes with local thermodynamic equilibrium. Examples include mechanical friction, fluid viscosity, electrical conduction / resistance, thermal conduction, and mass diffusion. Such models can be defined in terms of a function that maps thermodynamic forces, which drive the relaxation process, to thermodynamic fluxes, which seek to restore equilibrium. Based on the assumption that the microscopic dynamics are reversible, Onsager found that such maps must satisfy a symmetry property [15]. This motivates our definition of an Onsager structure.
Definition 6.12.
Let $I$ be an interface. An Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ is a fiber subbundle that admits the following representation:
$\mathcal{O}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,f\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,M(e)\,e\right\}\,,$
where $M$ is a smooth function that yields a symmetric non-negative definite linear operator $M(e)\colon\mathrm{T}^{*}\mathcal{X}_{I,p}\to\mathrm{T}\mathcal{X}_{I,p}$ for all effort variables $e\in\mathrm{T}^{*}\mathcal{X}_{I,p}$. ∎
In contrast to a Dirac structure, an Onsager structure is neither power-preserving, nor is it a locally linear relation of the port variables, since destruction of exergy, or equivalently production of entropy, is usually described by a non-linear relation. Hence, $\mathcal{O}\subset\mathcal{P}_{I}$ is not a vector subbundle of $\mathcal{P}_{I}$, but merely a fiber subbundle. The symmetry of $M(e)$ corresponds to Onsager’s reciprocal relations and its non-negative definiteness implies a non-negative exergy destruction rate, i.e. $\langle e\mid f\rangle\geq 0$.
Example 6.13.
Let the interface $I_{\text{mf}}=(N_{\text{mf}},\,\tau_{\text{mf}})$ be given by
$\begin{split}N_{\text{mf}}\>&=\>\{\mathtt{p},\,\mathtt{s}\}\\ \tau_{\text{mf}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\\ \tau_{\text{mf}}(\mathtt{s})\>&=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})\end{split}$
and let $d\in\mathbb{R}_{>0}$ be a friction coefficient. Then, the following defines an Onsager structure $\mathcal{O}_{\text{mf}}\subset\mathcal{P}_{I_{\text{mf}}}$ representing mechanical friction:
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}$
Here, $\upsilon=\mathtt{p.e}$ is the velocity and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s.e}$ is the absolute temperature at which kinetic energy is dissipated into the thermal energy domain, since adding ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels the shift $-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ in the effort variable, see 6.3. The phenomenological parameter $d$ could also be a function of the effort variables $e$, i.e. it could depend on the velocity and the temperature, or theoretically also the base point $x$. The above can be simplified to $\mathtt{p.f}=d\,\upsilon$ and $\mathtt{s.f}=-\frac{1}{\theta}\,d\,\upsilon^{2}$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels out in the flow variables, which here are the friction force and the entropy production rate (with a minus sign since thermal energy is leaving the component). The exergy destruction rate is given by $\langle\mathtt{p.e}\mid\mathtt{p.f}\rangle+\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\frac{d\,\upsilon^{2}}{\theta}\geq 0$, with $\frac{d\,\upsilon^{2}}{\theta}\geq 0$ being the entropy production rate. ∎
An irreversible component is defined by an Onsager structure that satisfies two additional conditions ensuring thermodynamic consistency.
Definition 6.14.
An irreversible component $(I,\,\mathcal{O})$ is given by an interface $I$ and an Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ such that energy and all quantities of the reference environment, except for entropy, are conserved and the time-reversal invariance property does strictly not hold. The condition for conservation of energy requires that $\forall\,e\in\mathrm{T}^{*}\mathcal{X}_{I,p}\,\colon$
$0\>=\>M(e)\,\bigl{(}\lambda(I)+e\bigr{)}\,.$
The condition that all quantities present in the reference environment, except for entropy, are conserved requires that $\forall\,p_{R}\in I_{R}$ with $\tau_{R}(p_{R})\neq((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$, we have that $\forall\,e\in\mathrm{T}^{*}\mathcal{X}_{I,p}\,\colon$
$0\>=\>M(e)\,\lambda(I,\,p_{R})\,.$
The time-reversal invariance property does strictly not hold if $\forall\,x\in\mathcal{X}_{I}\,\colon$
${(M(x))}^{ij}\neq 0\implies\mathrm{P}(x^{i})\>=\>\mathrm{P}({(M(x))}^{ij})\,\mathrm{P}(x^{j})\,.$
Here, we use the same notation as in 6.9 with indices $i$ and $j$ ranging over all power ports of $I$. ∎
From 6.4 and 6.3 we know that efforts variables are given by the differential of the energy function and additional shifts by the constant, intensive quantities of the environment, whenever the quantity of the respective port matches with a quantity present in the envirionment. We thus have $e=e^{\prime}-\lambda(I)$, where $e^{\prime}$ is precisely the part that stems from the differential of the energy. From this and the chain rule, it follows that energy is conserved if $0=\langle e^{\prime}\mid f\rangle=\langle e+\lambda(I)\mid M(e)\,e\rangle$. Taking the linear adjoint, this results in the condition in the definition, since $M(e)$ is symmetric. Regarding the condition that all quantities that are also present in the environment, except for entropy, are conserved, follows as already shown for reversible components, following 6.9. Again, regarding the time-reversal invariance condition, we refer to [46] or [3].
Example 6.15.
The irreversible component $(I_{\text{mf}},\,\mathcal{O}_{\text{mf}})$, as already defined in 6.13, models mechanical friction. The condition for conservation of energy is satisfied since
$\begin{bmatrix}0\\ 0\end{bmatrix}\>=\>d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\upsilon\\ \theta\end{bmatrix}\,.$
The condition that all quantities of the reference environment, except for entropy, are conserved is trivially satisfied. As required, the time-reversal invariance property does strictly not hold since
$\begin{split}\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,{(\mathrm{P}(\upsilon))}^{2}\,/\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\,.\end{split}$
Here, we used that temperature has positive parity and velocity has negative parity, which follows from the parities of entropy and momentum, knowing that energy has positive parity. ∎
Irreversible components, and hence EPHS in general, respect Onsager’s reciprocal relations because the symmetry of $M(e)$ in 6.12 implies equal cross-effect coefficients (see, e.g., [48]). In the terminology of Linear Irreversible Thermodynamics (LIT), thermodynamic fluxes are related to thermodynamic forces by a symmetric operator. The following example shows this in detail.
Example 6.16.
Let the interface $I_{\text{ht}}=(N_{\text{ht}},\,\tau_{\text{ht}})$ be given by
$\begin{split}N_{\text{ht}}\>&=\>\{\mathtt{s_{1}},\,\mathtt{s_{2}}\}\\ \tau_{\text{ht}}(\mathtt{s_{1}})\>&=\>\tau_{\text{ht}}(\mathtt{s_{2}})\>=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})\,,\end{split}$
let $\alpha\in\mathbb{R}_{>0}$ be a heat transfer coefficient, and let the Onsager structure $\mathcal{O}_{\text{ht}}\subset\mathcal{P}_{I_{\text{ht}}}$ be given by
$\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\alpha\,\begin{bmatrix}\frac{\theta_{2}}{\theta_{1}}&-1\\ -1&\frac{\theta_{1}}{\theta_{2}}\end{bmatrix}\,\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}\,,$
where $\theta_{1}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s_{1}.e}$ and $\theta_{2}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s_{2}.e}$ are the absolute temperatures of the two thermal energy domains. It can be easily checked that $(I_{\text{ht}},\,\mathcal{O}_{\text{ht}})$ defines an irreversible component. Again, the phenomenological parameter $\alpha$ could for instance be a function of the temperatures. The above can be simplified to $\mathtt{s_{1}.f}=-\frac{1}{\theta_{1}}\,\alpha\,(\theta_{2}-\theta_{1})$ and $\mathtt{s_{2}.f}=-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels out in the flow variables, which are rates of entropy change due to heat transfer. Since $M(e)$ is symmetric and non-negative definite, an Onsager structure can be represented in a factorized form:
$\underbrace{\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}}_{f}\>=\>\underbrace{\begin{bmatrix}+\frac{1}{\theta_{1}}\\ -\frac{1}{\theta_{3}}\end{bmatrix}}_{C(e)}\,\underbrace{\begin{bmatrix}\alpha\,\theta_{1}\,\theta_{3}\end{bmatrix}}_{D(e)}\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,\underbrace{\begin{bmatrix}+\frac{1}{\theta_{1}}&-\frac{1}{\theta_{3}}\end{bmatrix}}_{C^{*}(e)}\,\underbrace{\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}}_{e}$
The factorization has a physical interpretation in terms of LIT (see, e.g., [13]). The thermodynamic force that drives the irreversible process is given by
$X\>=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,C^{*}(e)\,e\>=\>\frac{1}{\theta_{2}}-\frac{1}{\theta_{1}}\,.$
Application of the symmetric operator $D(e)$ yields the thermodynamic flux
$J\>=\>D(e)\,\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,C^{*}(e)\,e\>=\>\alpha\,(\theta_{1}-\theta_{2})\,.$
Since $M(e)$ models a single process, $D(e)$ has rank $1$. Application of $C(e)$ yields the flow variables $f$. The net power $\langle\mathtt{s_{1}.e}\mid\mathtt{s_{1}.f}\rangle+\langle\mathtt{s_{2}.e}\mid\mathtt{s_{2}.f}\rangle=\langle e\mid f\rangle={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\frac{\alpha\,{(\theta_{1}-\theta_{2})}^{2}}{\theta_{1}\,\theta_{2}}\geq 0$ is the exergy destruction rate, and $\frac{\alpha\,{(\theta_{1}-\theta_{2})}^{2}}{\theta_{1}\,\theta_{2}}=\langle X\mid J\rangle\geq 0$ is the entropy production rate. ∎
The semantics of an irreversible component is simply given by its Onsager structure interpreted as a relation.
Definition 6.17.
The semantics assigned to irreversible component $(I,\,\mathcal{O})$ is the relation $\mathcal{O}:1\nrightarrow\mathcal{P}_{I}$ that is given by its Onsager structure $\mathcal{O}$. ∎
7 Composite systems
A composite system is formed by interconnecting a finite number of (primitive or composite) systems according to an interconnection pattern. The data defining a composite system consists of an interconnection pattern and, for each of its inner boxes, a system with the corresponding interface.
Definition 7.1.
Let $\mathsf{System}$ refer to the set of all systems defined with respect to a fixed reference environment. A composite system $(I,\,\gamma,\,J)$ is defined by the following data:
  1. 1.
    An interface $I$ of the composite system.
  2. 2.
    A directory of systems $\gamma\in\mathsf{Dtry}(\mathsf{System})$, which can naturally be reduced to a directory of interfaces $\beta\in\mathsf{Dtry}(\mathsf{Interface})$.
  3. 3.
    A partition $J$ of the combined interface $P=\sum\left[\mathtt{inner}\mapsto\sum\beta,\,\mathtt{outer}\mapsto I\right]$ such that $(\beta,\,I,\,J)$ defines an interconnection pattern. ∎
class of systems interface relation definition
[…]
Table 2: Overview of systems and their semantics given by a relation.
Table 2 recalls the different kinds of components defined in Section 6. In order to write the relations that determine their semantics in a unified form, the Cartesian unit is inserted a few times. As shown in the last row of the table, the semantics of a composite system takes the same form. We just need to compose the relation assigned to the interconnection pattern with the Cartesian product of the relations assigned to the subsystems filling its inner boxes.
Definition 7.2.
Let $(I,\,\gamma,\,J)$ be a composite system and let $P=(\beta,\,I,\,J)$ be its underlying interconnection pattern. Further, let $\mathcal{P}_{P}$ be the relation assigned to pattern $P$ according to 4.2. For any filled inner box $b\in\operatorname{dom}(\gamma)$, let $\left\llbracket{s}\right\rrbracket$ denote the relation assigned to the subsystem $s=\gamma(b)$. In case system $s$ is itself a composite system, the present definition applies recursively. Then, the semantics assigned to composite system $(I,\,\gamma,\,J)$ is the relation $\mathcal{C}:\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ given by
$\mathcal{C}\>=\>\mathcal{P}_{P}\circ\left(\prod_{b\,\in\,\operatorname{dom}(\gamma)}\left\llbracket{\gamma(b)}\right\rrbracket\right)\,.$
Here, $\mathcal{X}_{I_{s}}$ denotes the state space of all storage components and $\mathcal{X}_{I_{e}}$ denotes the state space of all environment components that are subsystems of the composite system in a recursive sense. ∎
Example 7.3.
Consider the composite system given by filling the inner boxes of the interconnection pattern for the nonisothermal damper shown in Figure 6 with three suitable systems. The system filling the box $\mathtt{tc}$ represents a thermal capacity and it is defined analogous to 6.3, but without a port for volume exchange. It gives a relation
$\mathcal{S}_{\text{tc}}:\mathrm{T}\mathcal{X}_{\text{tc}}\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{tc}}$
with $\mathcal{X}_{\text{tc}}=\mathbb{R}$. The system filling the box $\mathtt{mf}$ is a mechanical friction model as defined in 6.15 and it gives a relation
$\mathcal{O}_{\text{mf}}:1\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{mf}}$
with $\mathcal{X}_{\text{mf}}=\mathbb{R}\times\mathbb{R}$. The system filling the box $\mathtt{ht}$ is a heat transfer model as defined in 6.16 and it gives a relation
$\mathcal{O}_{\text{ht}}:1\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{ht}}$
with $\mathcal{X}_{\text{ht}}=\mathbb{R}\times\mathbb{R}$. Let the Cartesian product of the three relations be denoted by
$\mathcal{S}_{\text{tc}}\times\mathcal{O}_{\text{mf}}\times\mathcal{O}_{\text{ht}}\colon\mathrm{T}\mathcal{X}_{\text{tc}}\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{tc}}\times\mathbb{T}\mathcal{X}_{\text{mf}}\times\mathbb{T}\mathcal{X}_{\text{ht}}\,.$
Further, let the semantics of the interconnection pattern be given by the relation
$\mathcal{P}_{P}\colon\mathbb{T}\mathcal{X}_{\text{tc}}\times\mathbb{T}\mathcal{X}_{\text{mf}}\times\mathbb{T}\mathcal{X}_{\text{ht}}\nrightarrow\mathcal{P}_{I}\,,$
where $\mathcal{P}_{I}=\mathbb{T}\mathcal{X}_{\mathtt{p}}\times\mathbb{T}\mathcal{X}_{\mathtt{s}}=\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}$ is the bundle of port variables of the outer interface. Then, the semantics of the composite system is given by the relation
$\mathcal{C}\>=\>\mathcal{P}_{P}\circ\left(\mathcal{S}_{\text{tc}}\times\mathcal{O}_{\text{mf}}\times\mathcal{O}_{\text{ht}}\right)$
from $\mathrm{T}\mathcal{X}_{\text{tc}}\times 1$ to $\mathcal{P}_{I}$. ∎
For an isolated composite system the assigned relation $\mathcal{C}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow 1$ simply amounts to a subbundle of $\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}$, which implicitly defines a vector field on $\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$. Given an initial condition, one can integrate this vector field to determine the evolution of the system. We simply say that $(x,\,\dot{x})\in\mathcal{C}$, where $x=(x_{s},\,x_{e})$, gives the dynamics of the isolated system. The following definition generalizes this statement.
Definition 7.4.
Let $s\in\mathsf{System}$ be a system with interface $I$ and let the relation $\mathcal{R}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ define its semantics, according to the definitions mentioned in Table 2. Further, let $(x_{s},\,x_{e})\in\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$ denote the state variables associated to all (nested) storage and environment components. If the system is a reversible or an irreversible component or given by an interconnection of only such components this is vacuous. Moreover, let $(x,f,e)\in\mathcal{P}_{I}$ denote the port variables of its (outer) interface. If the system is isolated this is vacuous. Then, the dynamics of system $s$ is implicitly determined by
$\bigl{(}(x_{s},\,\dot{x}_{s}),\,(x_{e},\,\dot{x}_{e}),\,(x,\,f,\,e)\bigr{)}\in\mathcal{R}\,.\qed$
Example 7.5.
The dynamics of the composite system described in 7.3 is given by
$\bigl{(}(s,\dot{s}),\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e}),\,(\mathtt{s.x},\,\mathtt{s.f},\,\mathtt{s.e})\bigr{)}\in\mathcal{C}\,.$
This boils down to the following equations:
$\dot{s}\>$ $=\>\frac{1}{\theta_{1}}\left(d\,\upsilon^{2}\,+\,\alpha\,(\theta_{2}-\theta_{1})\right)$
$\mathtt{p.f}\>$ $=\>d\,\upsilon$
$\mathtt{s.f}\>$ $=\>-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$
[…]
In the context of programming languages, syntactically correct code does not necessarily define a semantically valid program. Similarly, there are cases where a compiler for the EPHS language must raise an error. For example, no junction can have two or more connected ports belonging to storage or environment components. A more comprehensive discussion of the conditions required for a composite system to be well-defined is left for future work.
8 Thermodynamic consistency
Akin to metriplectic systems or the GENERIC, the structure of EPHS ensures thermodynamic consistency. To simplify things a bit, we focus on isolated systems.
Proposition 8.1.
The dynamics of an isolated EPHS respect the first law and the second law. ∎
To prove this, one could first show that the semantics of any (isolated) composite system can be represented by some abstract system of evolution equations. Analogous to metriplectic systems or the GENERIC, one could then use the structural properties of this abstract system of equations to demonstrate thermodynamic consistency. The following proof, however, does not require this intermediate step.
Proof.
As the semantics associated to interconnection patterns is functorial, it suffices to show thermodynamic consistency for composite systems, where all subsystems are primitive. In terms of semantics, environment components are a special case of storage components. Any finite number of storage components can be combined into a single component by taking the named sum of their interfaces and by summing their energy functions. Similarly, any finite number of reversible/irreversible components can be combined into a single such component by taking the named sum of their interfaces and by taking the Cartesian product of the relations that define them. It hence suffices to show thermodynamic consistency for a composite system with one storage component, one reversible component, and one irreversible component.
  1. 1.
    First, we consider the storage component and focus specifically on the structure of its exergy function $H$, see 6.2. The following two facts follow immediately:
    1. (a)
      The energy $E$ must be conserved if the exergy $H$ and all quantities also present in the reference environment are conserved.
    2. (b)
      Entropy must be growing at a non-negative rate if the exergy $H$ is decreasing at a non-negative rate, while the energy $E$ and all quantities also present in the reference environment, except for entropy, are conserved.
  2. 2.
    Regarding the interconnection of the components based on 4.2, we know that at each junction the flow variables balance, while the effort variables are equal. We can conclude the following:
    1. (a)
      The interconnection is power-preserving, i.e. it conserves exergy.
    2. (b)
      The interconnection conserves the quantities associated with the ports because each flow variable represents the rate of change of the corresponding quantity, and the flow variables balance.
    3. (c)
      The interconnection conserves energy. This follows from the previous two points and point 1. (a)
  3. 3.
    Next, we consider the reversible component defined by a Dirac structure $\mathcal{D}$, see 6.7. Since $\mathcal{D}$ is a power-preserving relation, it conserves the exergy $H$. One of the conditions in 6.9 guarantees that $\mathcal{D}$ conserves all quantities also present in the reference environment, in particular entropy. Considering also point 1. (a), it follows that $\mathcal{D}$ must also conserve the energy $E$.
  4. 4.
    Finally, we consider the irreversible component defined by an Onsager structure $\mathcal{O}$, see 6.12. As a consequence of the non-negative definiteness property, $\mathcal{O}$ destroys exergy at a non-negative rate. One of the conditions in 6.14 guarantees that $\mathcal{O}$ conserves the energy $E$. Another condition guarantees that $\mathcal{O}$ conserves all quantities also present in the reference environment, except for entropy. Considering also point 1. (b), it follows that $\mathcal{O}$ must produce entropy at a non-negative rate.
The composite system must respect the first law, since the interconnection, the reversible component, and the irreversible component all conserve energy.
The composite system must respect the second law, since the interconnection and the reversible component conserve entropy, while the irreversible component produces entropy at a non-negative rate. ∎
The concept of a storage function was introduced in Willems’ theory of dissipative systems [49]. In this framework, EPHS without irreversible components are classified as lossless, whereas EPHS that include irreversible components are considered dissipative or passive. The additional conditions defining reversible and irreversible components in EPHS extend the notion of dissipativity, ensuring full thermodynamic consistency.
9 Example: shunt motor
This section compares a bond graph and an EPHS model of a direct current (DC) shunt motor. A DC motor has a fixed outer part, called stator, and a rotating inner part, called rotor. Both parts contain a coil of wire through which current flows. In the case of a shunt motor, these two inductors are connected in parallel to a source of electric energy. The working principle of such motors is based on a coupling of the electromagnetic and the kinetic energy domains via the Lorentz force and the commutation of the rotor coil after every half turn. However, both aspects are not directly resolved by the considered models. Instead, the coupling is considered to be lumped into a gyrator element.
Figure 10: Bond graph model of a DC shunt motor.
We start with the bond graph model that is taken from [8] and reproduced, with slightly different naming of variables, in Figure 10. The half arrows are called bonds and each bond has two associated variables, called flow and effort. The name of the flow variable is written on the side of the arrow head, while the name of the effort variable is written on the opposite side. However, not all variables are explicitly named in the example. At any 00-junction, the effort variables of all connected bonds are equal, while the flow variables balance, with their sign given by the bond direction. Dually, at $1$-junctions, flows are equal and efforts balance. Specifically, at the 00-junction on the left, the current from the $\mathrm{Se}$-element (effort source) with effort/voltage $E$ splits into the current $i_{s}$ through the stator coil and the current $i_{r}$ through the rotor coil. The inductance of the rotor coil is represented by the $\mathrm{I}$-element (generalized inductor) with inductance $L_{r}$, while an $\mathrm{I}$-element representing the stator coil inductance is omitted. The (steady-state) stator coil current $i_{s}$ is instead entirely determined by the $\mathrm{R}$-element (generalized resistor) at the top representing the coil’s electrical resistance $R_{s}$. The magnetic flux of the stator coil $b_{s}$ is then obtained as the signal that results from multiplying the signal $i_{s}$, which is the shared effort variable at the upper $1$-junction, with the constant gain $L_{s}$. The bond graph hence combines energy-based modeling with signal-based modeling based on block diagrams. At the $1$-junction on the left, the source voltage $E$ is equal to the sum of the voltage over the rotor coil inductance $L_{r}$, the voltage $u_{R}$ over the rotor coil resistance $R_{r}$, and the induced (back EMF) voltage $u_{r}$. The latter is determined by the $\mathrm{MGY}$-element (modulated gyrator), which describes the coupling of the electrical and the mechanical parts. The gyrator ratio is the signal $b_{s}$ coming from the gain block. At the gyrator, the effort variable of one port is equal to the product of the gyrator ratio and the flow variable of the other port. On the mechanical side, the load torque $M_{\text{load}}\leq 0$ is modeled by the effort source on the right, while the rotor’s angular mass $J_{r}$ is represented by an $\mathrm{I}$-element. At the $1$-junction the sum of the generated torque $M$ and the load torque $M_{\text{load}}$ is equal to the sum of the torque accelerating the angular mass and the friction torque $M_{R}$. The latter is the product of the friction coefficient $d_{r}$ and the angular velocity $\omega$. Putting all this together, the bond graph gives the system of ordinary differential equations
[…]
where
$i_{s}\>$ $=\>\frac{1}{R_{s}}\,E$
$b_{s}\>$ $=\>L_{s}\,i_{s}$
$u_{r}\>$ $=\>b_{s}\,\omega$
$M\>$ $=\>b_{s}\,i_{r}$
$M_{R}\>$ $=\>d_{r}\,\omega$
$u_{R}\>$ $=\>R_{r}\,i_{r}\,.$
As already shown in 4.3, the EPHS model is specified as an interconnection of two subsystems named $\mathtt{stator}$ and $\mathtt{rotor}$. The two systems have a common electric energy domain represented by the junction with connected ports named $\mathtt{q}$. The kinetic energy domain of the rotor is represented by the junction with connected ports named $\mathtt{p}$. Both energy domains are exposed such that the motor model can be integrated into another system. The conversion between electric and kinetic energy in the rotor depends on the state of the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{s}}$.
Figure 11: Interconnection pattern of the stator model.
The interconnection pattern of the stator model is shown in Figure 11. Box $\mathtt{emc}$ is filled by a reversible component that represents the coupling of the electric energy domain associated with port $\mathtt{q}$ and the magnetic energy domain associated with port $\mathtt{b_{s}}$. Box $\mathtt{coil}$ is filled by a storage component that represents the inductance of the stator coil. Further, box $\mathtt{res}$ is filled by an irreversible component that represents the resistance of the coil. Finally, box $\mathtt{env}$ is filled by an environment component whose thermal energy domain associated with port $\mathtt{s}$ absorbs the dissipated electromagnetic energy.
The reversible component $(I_{\text{emc}},\,\mathcal{D}_{\text{emc}})$ is defined by its interface
$I_{\text{emc}}=(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{emc}})$ with
$\tau_{\text{emc}}(\mathtt{q})\>$ $=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$
$\tau_{\text{emc}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
and its Dirac structure $\mathcal{D}_{\text{emc}}\subset\mathcal{P}_{I_{\text{emc}}}$ given by
$\begin{bmatrix}\mathtt{q.f}\\ \mathtt{b_{s}.f}\end{bmatrix}\>=\>\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\,\begin{bmatrix}\mathtt{q.e}\\ \mathtt{b_{s}.e}\end{bmatrix}\,.$
The storage component $(I_{\text{coil}},\,E_{\text{coil}})$ is defined by its interface $I_{\text{coil}}=(\{\mathtt{b_{s}}\},\,\tau_{\text{coil}})$ with $\tau_{\text{coil}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its energy function $E_{\text{coil}}:\mathcal{X}_{I_{\text{coil}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\mathtt{coil}}}=\mathbb{R}\ni\mathtt{b_{s}.x}=b_{s}$. The function is given by
$E_{\mathtt{coil}}(b_{s})\>=\>\frac{1}{2\,L_{s}}\,b_{s}^{2}\,,$
where the parameter $L_{s}$ is the inductance. The effort variable $\mathtt{b_{s}.e}=\frac{b_{s}}{L_{s}}$ is the current through the coil.
The irreversible component $(I_{\text{res}},\,\mathcal{O}_{\text{res}})$ is defined by its interface $I_{\text{res}}=(\{\mathtt{b_{s}},\,\mathtt{s}\},\,\tau_{\text{res}})$ with
$\tau_{\text{res}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
$\tau_{\text{res}}(\mathtt{s})\>$ $=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$
and its Onsager structure $\mathcal{O}_{\text{res}}\subset\mathcal{P}_{\text{res}}$ given by
$\begin{bmatrix}\mathtt{b_{s}.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,R_{s}\,\begin{bmatrix}\theta&-i\\ -i&\frac{i^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{b_{s}.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}R_{s}\,i\\ -\frac{1}{\theta}\,R_{s}\,i^{2}\end{bmatrix}\,,$
where the parameter $R_{s}$ is the resistance of the coil, $i=\mathtt{b_{s}.e}$ is the current through the coil, and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}+}\mathtt{s.e}$ is the absolute temperature at which heat is dissipated into the thermal energy domain.
The environment component is defined by the subinterface $I_{\text{env}}=\{\mathtt{s}\}\subseteq I_{R}$.
Figure 12: Interconnection pattern of the rotor model.
The interconnection pattern of the rotor model is shown in Figure 12. The components filling the boxes $\mathtt{coil}$, $\mathtt{emc}$, $\mathtt{res}$ and $\mathtt{env}$ are defined as for the stator model, except for a different port name, namely $\mathtt{b}$ instead of $\mathtt{b_{s}}$, and different parameters, namely $L_{r}$ instead of $L_{s}$ and $R_{r}$ instead of $R_{s}$. Further, box $\mathtt{mkc}$ is filled by a reversible component that describes the coupling of the magnetic energy domain associated with port $\mathtt{b}$ and the (rotational) kinetic energy domain associated with port $\mathtt{p}$. This coupling depends on the magnetic flux $\mathtt{b_{s}.x}$ that is generated by the stator coil. Box $\mathtt{mass}$ is filled by a storage component that represents the rotor’s angular mass and box $\mathtt{mf}$ is filled by an irreversible component that represents its mechanical friction.
The reversible component $(I_{\text{mkc}},\mathcal{D}_{\text{mkc}})$ is defined by its interface
$I_{\text{mkc}}=(\{\mathtt{b},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{mkc}})$ with
$\tau_{\text{mkc}}(\mathtt{b})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
$\tau_{\text{mkc}}(\mathtt{p})\>$ $=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
$\tau_{\text{mkc}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
and its Dirac structure $\mathcal{D}_{\mathtt{mkc}}\subset\mathcal{P}_{I_{\text{mkc}}}$ given by
$\begin{bmatrix}\mathtt{b.f}\\ \mathtt{p.f}\end{bmatrix}\>=\>\begin{bmatrix}0&\mathtt{b_{s}.x}\\ -\mathtt{b_{s}.x}&0\end{bmatrix}\,\begin{bmatrix}\mathtt{b.e}\\ \mathtt{p.e}\end{bmatrix}\,.$
The storage component $(I_{\text{mass}},\,E_{\text{mass}})$ is defined by its interface $I_{\text{mass}}=(\{\mathtt{p}\},\,\tau_{\text{mass}})$ with $\tau_{\text{mass}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$ and its energy function
$E_{\text{mass}}:\mathcal{X}_{I_{\text{mass}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\text{mass}}}=\mathbb{R}\ni\mathtt{p.x}=p$. The function is given by
$E_{\text{mass}}(p)\>=\>\frac{1}{2\,J_{r}}\,p^{2}\,,$
where the parameter $J_{r}$ is the angular mass.
The irreversible component $(I_{\text{mf}},\mathcal{O}_{\text{mf}})$ is defined by its interface $I_{\text{mf}}=(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{mf}})$ with
$\tau_{\text{mf}}(\mathtt{p})\>$ $=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
$\tau_{\text{mf}}(\mathtt{s})\>$ $=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$
and its Onsager structure $\mathcal{O}_{\text{mf}}\subset\mathcal{P}_{I_{\text{mf}}}$ given by
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,d_{r}\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}d_{r}\,\upsilon\\ -\frac{d_{r}\,\upsilon^{2}}{\theta}\end{bmatrix}\,,$
where the parameter $d_{r}$ is the mechanical friction coefficient, $\upsilon=\mathtt{p.e}$ is the angular velocity, and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}+}\mathtt{s.e}$ is the absolute temperature at which heat is dissipated into the thermal energy domain.
We write $(b_{s},\,b,\,p,\,s)\in\mathbb{R}^{4}$ for the state of the interconnected motor model, where we identify the entropy state variables $s$ of both environment components. In a future work, the identification of environment components can be formalized by composing with a suitable relation. The EPHS model then yields the equations
$\dot{b}_{s}\>$ $=\>\mathtt{q.e}-R_{s}\,i_{s}$
$\dot{b}\>$ $=\>\mathtt{q.e}-R_{r}\,i_{r}-b_{s}\,\omega$
$\dot{p}\>$ $=\>b_{s}\,i_{r}-d_{r}\,\omega+\mathtt{p.f}$
$\dot{s}\>$ $=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\left(R_{s}\,i_{s}^{2}+R_{r}\,i_{r}^{2}+d_{r}\,\omega^{2}\right)$
$\mathtt{q.f}\>$ $=\>i_{s}+i_{r}$
$\mathtt{p.e}\>$ $=\>\omega\,,$
where
$i_{s}\>$ $=\>\frac{1}{L_{s}}\,b_{s}$
$i_{r}\>$ $=\>\frac{1}{L_{r}}\,b$
$\omega\>$ $=\>\frac{1}{J_{r}}\,p\,,$
which are similar to the equations for the bond graph model, except that (1) the source of electrical energy and the mechanical load are regarded as external dynamical systems, rather than internal constants, and (2) no quasi-static assumption is hence made for the magnetic field of the stator.
10 Discussion
10.1 Comparison with bond graphs
Bond graphs provide a graphical notation for representing physical systems in terms of subsystems that exchange energy. However, the notation does not naturally support the hierarchical nesting of systems. Relying exclusively on primitive subsystems limits the applicability of bond graphs to complex systems, where a hierarchical breakdown into simpler parts would facilitate both model construction and reuse.
Although the notation is not inherently composable, one can introduce conventions that allow for subsystems defined by open bond graphs. For example, The convention in [50] considers bond graphs that share a specific element in common as a modules. These modules can be ‘glued’ together in a parent-level bond graph via an additional 00-junction. In this approach, the common element in each individual bond graph is replaced by an external port; the element is instantiated once in a parent-level bond graph, where it is connected to a 00-junction. The modules are represented by additional elements, and their ports are also connected to the 00-junction. The direction of these bonds must match the direction of the corresponding bond (or port) at the module level.
While port-Hamiltonian systems can be seen as a geometric formalization of the dynamical systems derived from generalized bond graphs, the algebraic formalization of the bond graph notation itself is not straightforward (see, e.g., [51, 32]). Some inherent complexity in the notation is evident also in practice, where the direction in which exchange of power is counted as positive, and sometimes also the role of a variable as flow versus effort, must be decided on a bond-by-bond basis. These decisions must either be left to the user or managed through heuristic algorithms or conventions. This can impede seamless hierarchical decomposition and reuse of subsystems.
Interconnection patterns provide a composable, graphical syntax for the EPHS modeling language. Compared to bond graphs, interconnection patterns have an outer box, and they exclusively rely on 00-junctions. Based on this simplification, the mentioned issues are resolved:
  • 1.
    Interconnection patterns have a clear mathematical structure, which supports their hierarchical (de-)composition.
  • 2.
    Interconnection patterns are relatively easy to interpret, owing to their simple structure, the direct correspondence between junctions and energy domains, and the clear role of a power variables as either flow or effort.
  • 3.
    The direction in which exchange of power is counted positive is predetermined in agreement with the convention used in thermodynamics, i.e. energy supplied to a system has a positive sign.
In contrast to bond graphs, EPHS models are guaranteed to be thermodynamically consistent.
Bond graphs may incorporate elements from block diagrams. In addition to bonds, which represent energy exchange, there may be edges that transmit signals according to their orientation. Elements may have signal ports to receive signals, which modulate their behavior, and there may be ‘blocks’ that perform arbitrary signal processing. In contrast, the formalism presented here does not cover arbitrary signal processing; system behavior can be modulated only by state variables of other systems. To express this, every port has a state variable, which is shared with all other ports connected to the same junction. We distinguish between state ports, which share only state, and power ports, which additionally model energy exchange.
10.2 Comparison with port-Hamiltonian systems
Port-Hamiltonian Systems (PHS) provide a mathematical framework for open dynamical systems with guaranteed passivity. While bond graphs are frequently used to depict port-Hamiltonian models, establishing a rigorous link between bond graph syntax and port-Hamiltonian semantics remains challenging. On the one hand, bond graphs offer a network representation of individual systems, but the representation is not inherently composable. On the other hand, PHS provide a composable representation of individual systems, albeit without an explicit internal network structure. In particular, the PHS framework provides no built-in support for referencing subsystems and their ports. Previous work has focused on algorithms for converting certain classes of bond graphs to PHS (see, e.g., [52]).
A well-known issue concerns the relationship between port-Hamiltonian structure, which implies passivity, and its physical interpretation. Owing to its roots in mechanics, the ‘Hamiltonian’ storage function is typically interpreted as the system’s energy. However, the differential of the ‘Hamiltonian’ generally also drives dissipative dynamics, resulting in an apparent loss of ‘energy’. This raises questions, especially outside the realm of electromechanical systems, where thermodynamic considerations are often overlooked. In response, some members of the port-Hamiltonian community have proposed alternative port-based frameworks that explicitly encode thermodynamic structure (see, e.g., [53, 54, 55]).
EPHS provide a mathematical framework for open dynamical systems with guaranteed thermodynamic consistency. In contrast to PHS, this framework is directly based on a composable, graphical syntax used to specify how more complex systems are formed by interconnecting simpler systems. An explicit understanding of the translation between graphical syntax and relational semantics guarantees that syntax and semantics compose in compatible ways. This enables a modular and hierarchical modeling approach, which inherently supports referencing nested subsystems and their ports. In a sense, this approach unites the strengths of both bond graphs and PHS.
The dynamical systems defined by EPHS models essentially possess port-Hamiltonian structure. Beyond passivity, additional structural properties imposed on the primitive systems guarantee that all EPHS share a uniform and consistent physical interpretation, aligned with nonequilibrium thermodynamics. Since EPHS compose simply by sharing energy domains, users typically do not need to explicitly define external ports. As a result, users can typically integrate existing models into more complex systems without any prior adaptation.
10.3 Comparison with metriplectic and GENERIC systems
Metriplectic systems and the GENERIC provide a framework for (usually isolated) dynamical systems with guaranteed thermodynamic consistency. The dynamical systems defined by (isolated) EPHS models are basically metriplectic systems or instances of the GENERIC, since the built-in thermodynamic consistency follows from essentially the same ideas. Regarding the reversible part, EPHS appear to be more general, since they use Dirac structures, rather than the less general Poisson structures used in the definition of metriplectic systems or the GENERIC. This allows EPHS to easily model kinematic constraints, which is important e.g. for multibody systems (see [56]). However, it should be no problem to generalize the metriplectic or GENERIC framework accordingly. Regarding the irreversible part, EPHS appears to be less general, since it remains unclear if the single-generator formulation based on exergy is useful beyond local thermodynamic equilibrium.
In contrast to the metriplectic or GENERIC framework, EPHS enable a modular and hierarchical approach to thermodynamic modeling, based on a graphical syntax for building more complex models from simpler parts.
10.4 Future work
Future work shall address the details omitted in this work. A more technical paper will define interconnection patterns and their functorial semantics using a categorical framework based on directories and the theory of operads/multicategories.
As presented here, the definition of (composite) EPHS exhibits a recursive structure. It would be interesting to study the ‘closure’ of all EPHS models by deriving representations of EPHS as structured systems of differential-algebraic equations. The analysis could benefit further developments and further clarify the relationship with the metriplectic or GENERIC framework as well as port-Hamiltonian systems.
As an important groundwork for further practical developments, we have implemented the framework within the Julia programming language, see [57]. The reader is invited to review its application in several examples. Based on this, future work can explore topics such as advanced applications (see, e.g., [56, 11]), structure-preserving numerical methods, semi-automated control design, and the integration with scientific machine learning methods.
Author contribution statement
Markus Lohmayer: Conceptualization, Investigation, Writing – Original Draft, Writing – Review & Editing, Visualization; Owen Lynch: Investigation, Writing – Original Draft; Sigrid Leyendecker: Supervision, Review & Editing, Funding
Acknowledgments
We thank David Spivak for help when working out the concept of namespaces. The second author acknowledges support from the DARPA ASKEM program under grant HR00112220038.